### v1

# Scaling RL to Long Videos

Abstract
Abstract: We introduce a full-stack framework that scales up reasoning in vision-language models (VLMs) to long videos, leveraging reinforcement learning. We address the unique challenges of long video reasoning by integrating three critical components: (1) a large-scale dataset, LongVideo-Reason, comprising 52K long video QA pairs with high-quality reasoning annotations across diverse domains such as sports, games, and vlogs; (2) a two-stage training pipeline that extends VLMs with chain-of-thought supervised fine-tuning (CoT-SFT) and reinforcement learning (RL); and (3) a training infrastructure for long video RL, named Multi-modal Reinforcement Sequence Parallelism (MR-SP), which incorporates sequence parallelism and a vLLM-based engine tailored for long video, using cached video embeddings for efficient rollout and prefilling. In experiments, LongVILA-R1-7B achieves strong performance on long video QA benchmarks such as VideoMME. It also outperforms Video-R1-7B and even matches Gemini-1.5-Pro across temporal reasoning, goal and purpose reasoning, spatial reasoning, and plot reasoning on our LongVideo-Reason-eval benchmark. Notably, our MR-SP system achieves up to 2.1$\times$ speedup on long video RL training. LongVILA-R1 demonstrates consistent performance gains as the number of input video frames scales. LongVILA-R1 marks a firm step towards long video reasoning in VLMs. In addition, we release our training system for public availability that supports RL training on various modalities (video, text, and audio), various models (VILA and Qwen series), and even image and video generation models. On a single A100 node (8 GPUs), it supports RL training on hour-long videos (e.g., 3,600 frames / around 256k tokens). Code and models are available at https://github.com/NVlabs/Long-RL.
1 Introduction
Understanding long videos requires more than simple recognition—it demands reasoning from temporal, spatial, goal-oriented, and narrative perspectives [1]. As illustrated in Figure 2, answering high-level questions often hinges on a model’s ability to integrate clues distributed across time, infer hidden goals or strategies, track entities spatially, and comprehend the evolving plot. For instance, predicting the winner of a football penalty shootout involves assessing emotional cues and tactical behavior (temporal and goal reasoning), while determining the final location of a hidden ball requires precise spatial tracking. Likewise, evaluating a poker player’s decision demands interpreting implicit strategies beyond surface actions (goal reasoning) and understanding a character’s development or match trajectory reflects the need for plot reasoning. These examples underscore that reasoning is indispensable for long video understanding that goes beyond recognition alone.
Figure 1: Training efficiency comparison with MR-SP (SP degree=4) on Qwen2.5-VL-7B and LongVILA-R1-7B and a single node 8$\times$ A100 GPUs. It achieves 2.1$\times$ speed-up and avoids GPU OOM issue on long frames.
Figure 2: Examples of LongVILA-R1. The illustration demonstrates sample tasks and their reasoning. From left to right, the examples include predicting the results of a football match, decision-making reasoning in Texas Hold’em Poker, and trajectory for spatial dynamics of objects. Notably, the spatial tracking video involves a relatively complex dynamic moving, for which the model fails to achieve accurate reasoning until the number of input video frames increases to 128.
Despite the clear importance of reasoning in long video understanding, enabling such capabilities in long-video VLMs poses significant challenges [2, 3, 4, 5, 6]. First, the collection of high-quality long video reasoning datasets is inherently difficult. Unlike domains such as math or code reasoning, where structured supervision and benchmarks are readily available [7, 8], long video reasoning requires annotating complex temporal dynamics, goals, spatial relations, and narrative elements—often across minutes or hours of footage [9]. This process is labor-intensive and subjective, making large-scale dataset construction slow and costly. Second, the RL training framework for long videos is challenging. Reinforcement learning, a common strategy for aligning models with complex reasoning objectives, is computationally expensive and sample-inefficient [10]. When applied to long videos, RL becomes even more burdensome due to the extended video frames, requiring more memory and longer rollout runtime. These challenges jointly hinder the development of effective long video VLMs with strong reasoning capabilities.
In this work, we introduce LongVILA-R1, a comprehensive framework exploring the reasoning capabilities for long-video understandings. Firstly, we strategically construct a high-quality dataset with CoT annotations for long video reasoning, named LongVideo-Reason. Leveraging a powerful VLM (NVILA-8B) [11] and a leading open-source reasoning LLM, we develop a dataset comprising 52K high-quality Question-Reasoning-Answer pairs for long videos. We use 18K high-quality samples for Long-CoT-SFT to initialize the model’s reasoning and instruction-following abilities, and 33K samples with an additional 110K video data [2] for reinforcement learning. This two-stage training combines high-quality reasoning annotations with reinforcement learning, enabling LongVILA-R1 to achieve superior and generalized video reasoning. We also manually curate a balanced set of 1K long-video samples to build a new benchmark, LongVideo-Reason-eval, that evaluates performance from four perspectives: Temporal, Goal and Purpose, Spatial, and Plot and Narrative, for a comprehensive assessment.
Figure 3: Data Distribution of LongVideo-Reason and total data in the LongVILA-R1 training framework. LongVideo-Reason comprises a total of 18,077 videos, from which 52K QAs with reasoning annotations. Additionally, we also include additional 110K QAs from existing works [12, 13, 14, 15, 16] for RL training.
Secondly, we propose a training framework for VLMs to advance long video reasoning. As illustrated in Figure 5, this framework incorporates two stages, i.e., Stage-1: Long CoT-SFT, and Stage-2: RL for long video reasoning. To address the unique challenges of long video RL, including massive visual embeddings, heavy rollouts, and long-context LLM prefilling, we develop an efficient and scalable solution, referred to as Multi-modal Reinforcement Sequence Parallelism (MR-SP). It incorporates a vLLM engine [17] tailored for LongVILA and a caching scheme for video embeddings. The MR-SP system alleviates the problem of intensive memory and facilitates RL training of long video VLMs. As shown in Figure 1, our MR-SP system achieves up to 2.1$\times$ speedup on 512-frame video RL training on 7B models and enables longer training frames without out-of-memory (OOM).
LongVILA-R1-7B shows strong performance on general video benchmark, e.g., 68.4% on VideoMME [18](with subtitle). On our LongVideo-Reason-eval benchmark, LongVILA-R1-7B achieves an average accuracy of 67.9% across four reasoning categories, surpassing open-source models including Video-R1-7B [2] and proprietary models such as GPT-4o [19] by a large margin, and matches the performance of Gemini-1.5-Pro [20]. In addition, LongVILA-R1 exhibits progressive performance improvements as the number of input frames increased.
2 Related Work
Multi-modal reasoning models.
The field of multi-modal reasoning has advanced significantly, particularly in Vision-Language Models (VLMs). GPT-4o [19] improves visual understanding through enhanced reasoning, while Gemini-1.5-Pro [20] extends context length to 1 million tokens, achieving state-of-the-art performance in VideoMME [18]. Following the substantial progress of architecture and training algorithms in open-source VLMs [21, 11, 22, 23, 24], multi-modal reasoning has been further explored in works including LMM-R1 [25] which employs a two-stage training strategy, Vision-R1 [26] that addresses post-cold-start overthinking, and Video-R1 [2] which enhances RL for video via T-GRPO in 16 frames. However, these approaches primarily focus on single images or short videos, and long video reasoning still poses great challenges.
Sequence parallelism.
Training with long contexts often exceeds the memory capacity of a single device [27], necessitating efficient distribution strategies. Sequence parallelism (SP) has become a widely adopted solution  [28, 29, 30, 31, 32]. For example, ring-based systems like LightSeq [33] and Ring Attention [28] use point-to-point (P2P) communication, while DeepSpeed-Ulysses [29] employs all-to-all (A2A) primitives to optimize attention computations. Additionally, USP and LoongTrain  [31, 32] were introduced to integrate Ring-style SP and Ulysses SP. LongVILA [4] further proposed multi-modal SP (MM-SP), enabling vision-language models to handle long-context inputs. However, multi-modal reinforcement learning introduces additional challenges, as it requires extensive sampling from long, mixed-token sequences [10], particularly in complex group optimization tasks [7].
RL frameworks for LLMs/VLMs.
Reinforcement Learning (RL) has become a key strategy for enhancing Large Language Models (LLMs), particularly through Reinforcement Learning with Human Feedback (RLHF) [34] or Direct Preference Optimization [35], which aligns model outputs with human preferences. Recent advancements demonstrate that RL significantly improves LLM reasoning abilities. For example, DeepSeek-R1 [36] utilizes the Group Relative Policy Optimization (GRPO) algorithm [37], integrating group-based sampling and rule-based rewards. On the other hand, RL [38] poses a unique challenge of heavy computational cost, especially in multi-modal settings. To address this problem, HybridFlow [10] is introduced, leveraging Ray [39] for efficient data flow and vLLM [17] for faster sampling. Nevertheless, this remains a bottleneck when processing long video sequences, with group-based sampling constrained by the high computational cost of long-context sampling and visual encoding. In this work, we propose MR-SP that provides up to 2.1$\times$ speedup for long video RL training.
Figure 4: Data generation process for the LongVideo-Reason dataset. This process begins with segmenting videos into 10-second clips and generating captions for each clip using NVILA-8B. Then based on the captions of all clips in a video, we generate question-answer pairs that involve reasoning across the content of the whole video, along with the reasoning annotations using a leading open-source reasoning LLM. Reasoning questions are categorized into Temporal, Goal and Purpose, Spatial, and Plot and Narrative. Finally, the reasoning annotations are reformatted for conciseness and alignment with video details. We present a more detailed figure of data generation process in Figure 9 in the appendix.
3 LongVideo-Reason Data Construction
High-quality annotated datasets for long videos are critical for VLMs. Existing publicly available long-video datasets lack high-quality reasoning annotations. In this section, we detail the data construction process (Figure 4) of a 52K long-video QAs with reasons.
3.1 Overview of Data Curation
We first curate 18K long videos from the Shot2Story dataset [40] (Figure 3, left). We then apply a high-quality automated annotation pipeline for CoT as detailed in Section 3.2, and end up with a total of 52K Question-Reasoning-Answer pairs where each sample, based on the type of question it is reasoning about, can be categorized into Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning (Figure 3, middle). This dataset is designed to support various types of long-video reasoning tasks comprehensively.
Given the sensitivity of GRPO to batch sampling [41, 25], a data filtering approach is adopted. Specifically, we employ a test-scaling method where LongVILA [4] performs inference 10 times on the original datasets. Questions consistently answered correctly or incorrectly are labeled as "easy" or "hard" while those inducing diverse predictions are labeled as "medium". We use both the "easy" and "hard" samples for Stage-1 COT-SFT and use the "medium" samples for Stage-2 RL training. The reason is that GRPO expects different rollouts of each sample to be diverse in order to have meaningful advantages, and the gradient vanishes if all the rollouts predict correct or incorrect answers. The COT-SFT subset (18K) features high-quality CoT reasoning processes formatted in a standard <think></think><answer></answer> structure, providing abundant resources for warm-up training during Stage 1 of the model’s reasoning capabilities. Meanwhile, the RL subset contains 33K challenging long-video Q&A, which is leveraged in Stage 2 for scaling reasoning through reinforcement learning. To further enhance RL scaling, we incorporate an additional 110K high-quality open-source videos (Figure 3, right) from other datasets [12, 13, 14, 15, 16]. This combination improves the model’s generalization.
Figure 5: The LongVILA-R1 training pipeline. LongVILA-R1 builds upon the base training pipeline for LongVILA. MM-SP is further employed for SFT on long video understanding tasks with long CoT. Then, reinforcement scaling learning is conducted through Multi-modal Reinforcement Sequential Parallelism (MR-SP).
Figure 6: LongVILA-R1 RL training framework. The framework integrates multi-modal reinforcement sequence parallelism (MR-SP) for scalable video frame encoding and LLM prefilling. RL utilizes a vLLM-based engine with cached video embeddings, tailored for LongVILA rollout. Rewards for accuracy and format guide policy optimization.
3.2 Long-Video Reasoning Generation
We introduce an automated annotation pipeline (Figure 4) that generates high-quality Question-Reasoning-Answer pairs from long videos. This process begins by segmenting videos into short clips (~10 seconds each), each of which is annotated using the NVILA-8B [11] model to provide descriptive captions. Leveraging the breakthrough in text-based reasoning, we then deploy a leading open-source reasoning LLM, provide the captions of all the clips in each video, and prompt it to generate diverse types of Question-Reasoning-Answer pairs that involve reasoning over the content across the whole video. Specifically, we design four types of prompts to encourage the LLM to generate a Question-Reasoning-Answer pair that focuses on one of the four types of reasoning: Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning. To ensure VLMs focus on visual details, we also craft the prompts with phrases such as “checking the video” and “analyzing the scene”, which guide iterative examination of visual content. Finally, an LLM is then used to refine and streamline the reasoning steps. We also manually curated 1,000 high-quality complex reasoning questions across four reasoning categories to serve as a new benchmark (LongVideo-Reason-eval) for evaluating VLMs in reasoning abilities. This entire data procedure consumes about 40,000 GPU hours on H100s.
4 LongVILA Training Pipeline
As shown in Figure 5, there are two extended training stages in LongVILA-R1, i.e., (1) warm-up for long video reasoning, utilizing 18K data with high-quality CoT for SFT on the MM-SP system [4]; (2) reinforcement learning with dense frames from long videos.
4.1 Long Video CoT Supervised Fine-Tuning
Utilizing 52K high-quality question-reasoning-answer pairs, we apply the data filtering method described in Section 3.1 to select 18K examples for long CoT-SFT, serving as a warm-up phase for subsequent RL. This stage equips the model with fundamental reasoning abilities and instruction-following skills for long video scenarios. To efficiently perform SFT on hundreds of frames, we adopt the MM-SP [4] training system from LongVILA. As demonstrated in Section 6.2, SFT solely on our LongVideo-Reason dataset also effectively improves the model with foundational reasoning capabilities.
Figure 7: The workflow of multi-modal reinforcement sequential parallel (MR-SP). To accommodate multi-modal input in reinforcement learning, we develop a custom sharding strategy to ensure balanced workload distribution and compatibility with SP communication. Efficient video embedding reuse and vLLM rollout acceleration strategies are implemented, while meeting the demands of policy model prefilling for dense video frames.
4.2 GRPO for Long Video
Building on the advancements of the GRPO [7] algorithm and prior explorations of multi-modal reasoning training [2, 25], we adhere to the standard GRPO framework to train our model. For each given question $q$, the policy model generates a group of candidate responses $\{o_{1},o_{2},...,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$, accompanied by their corresponding rewards $\{r_{1},r_{2},...,r_{G}\}$, which are computed based on rule-based reward functions (format/accuracy). The model $\pi_{\theta}$ is subsequently optimized by maximizing the following objective function:
$\mathcal{J}(\theta)=\mathbb{E}_{q,\{o_{i}\}}[\frac{1}{G}\sum_{i=1}^{G}(\mathrm{min}(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}A_{i},\\ \mathrm{clip}(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)},1-\epsilon,1+\epsilon)A_{i})-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{ref}))]$ (1)
where $\epsilon$ and $\beta$ are hyper-parameters, $G$ is set as 8 in our experiments, and the sampled rewards above are normalized to get the advantages ($A_{i}$) for updating the model:
$A_{i}=\frac{r_{i}-\mathrm{mean}(\{r_{1},r_{2},...,r_{G}\})}{\mathrm{std}(\{r_{1},r_{2},...,r_{G}\})}$ (2)
However, RL for long videos presents significant challenges due to the high computational demands of processing hundreds to thousands of frames. Existing RL frameworks struggle with such long-context training in rollout and LLM prefilling. To address this, we develop the MR-SP framework (Section 5), which efficiently scales reinforcement learning for long-context video reasoning.
Considering the sensitivity of GRPO to sampling during training [41], we use the 33K filtered data for reinforcement learning as described in Section 3.1. Additionally, an extra 110K samples from Video-R1 [2] are incorporated to scale up the RL. This approach aims to guide the model in freely exploring and developing more effective and generalized reasoning strategies.
5 Multi-modal Reinforcement SP
Existing RL frameworks for VLMs, such as R1-V [38] and EasyR1 [42], are not designed for long videos which present unique challenges due to their high token volume. To address this, we introduce Multi-modal Reinforcement Sequence Parallelism (MR-SP), a framework for efficient RL training on long videos. MR-SP leverages sequence parallelism in both rollout and pre-filling stages, enabling long videos in RL, with reduced overhead.
Table 1: Performance on 9 video benchmarks [43, 44, 45, 46, 14, 47, 13, 48, 18]. LongVILA-R1-7B outperforms LongVILA-7B across all benchmarks, with varying margins that reflect different levels of reasoning difficulty.
Model
ActNet-QA
EgoSchema
EventBench
LVideoBench
PercepTest
MVBench
NExT-QA
VNBench
VideoMME
test test val val val test mc val w/o sub. w/ sub.
GPT-4V [49] 57.0 - 32.6 61.3 - 43.5 - - 59.9 63.3
GPT-4o [19] - - 53.3 66.7 - - - 64.4 71.9 77.2
Gemini-1.5-Pro [20] 57.5 72.2 43.2 64.0 - - - 66.7 75.0 81.3
Video-LLaVA-7B [50] 45.3 38.4 5.9 37.6 - 43.5 - 12.4 39.9 41.6
Flash-VStream-7B [51] 51.9 - - - - - 61.6 - - -
ShareGPT4Video-8B [52] 50.8 - - 41.8 - 51.2 - - 39.9 43.6
VideoLLaMA2-7B [53] 50.2 51.7 6.9 - 51.4 54.6 - 4.5 47.9 50.3
VideoLLaMA2.1-7B [53] 53.0 53.1 - - 54.9 57.3 - - 54.9 56.4
Kangaroo-8B [54] - 62.7 - 54.8 - 61.1 - - 56.0 57.6
PLLaVA-7B [55] 56.3 - 28.2 39.2 - 46.6 - - - -
LLaVA-OV-7B [56] 56.7 60.1 - 56.4 57.1 56.7 79.4 51.8 58.2 61.5
LongVILA-7B [4] 59.5 67.7 58.0 57.1 58.1 67.1 80.7 63.0 60.1 65.1
LongVILA-R1-7B 59.6 68.1 58.2 57.6 59.3 67.6 80.8 66.8 62.4 68.4
Table 2: Performance on LongVideo-Reason-eval. LongVILA-R1-7B achieves the highest overall score.
     Model      Temporal      Goal      Plot      Spatial      Overall
     GPT-4o [19]      64.4      62.4      60.4      55.6      60.7
     Gemini-1.5-Pro [20]      73.2      66.4      66.4      63.2      67.3
     Video-R1-7B [2]      66.4      62.8      63.6      58.0      62.7
     LongVILA-R1-7B      71.6      66.4      63.6      70.0      67.9
Table 3: Performance comparison on VideoMME [18] benchmark in details.
Model w/o subtitle w subtitle
Overall Short Medium Long Overall Short Medium Long
Video-LLaVA-7B [50] 39.9 45.3 38.0 36.2 41.6 46.1 40.7 38.1
SliME-8B [57] 45.3 53.3 55.4 39.8 47.2 55.4 44.4 41.7
ShareGPT4Video-8B [52] 39.9 48.3 36.3 35.0 43.6 53.6 39.3 37.9
VideoChat2-7B [58] 39.5 48.3 37.0 33.2 43.8 52.8 39.4 39.2
VideoLLaMA2-7B [53] 47.9 56.0 45.4 42.1 50.3 59.4 47.6 43.8
Chat-Univi-v1.5-7B [59] 40.6 45.7 40.3 35.8 45.9 51.2 44.6 41.8
Kangaroo-8B [54] 56.0 66.1 55.3 46.7 57.6 68.0 55.4 49.3
ShareGemini-7B [60] 43.2 49.1 41.3 39.1 47.9 49.1 47.3 43.4
LongVA-7B [1] 52.6 61.1 50.4 46.2 54.3 61.1 53.6 47.6
Video-R1-7B [2] 59.3 - - - - - - -
LongVILA-7B [4] 60.1 69.0 58.3 53.0 65.1 72.9 64.9 57.4
LongVILA-R1-7B 62.4 71.9 62.0 53.3 68.4 76.2 69.8 59.1
5.1 Stage 1 - Rollout with Paralleled Encoding
To support long-video reinforcement learning efficiently, we adopt sequence parallelism (SP) for the video encoding stage. As shown in Figure 7, the input video frames are first evenly divided across multiple GPUs (e.g., GPU 1 to GPU 3), each equipped with its own vision tower. Each GPU independently processes a slice of the video, encoding only a subset of the frames. The resulting video embeddings are then aggregated with text embeddings via an all-gather operation as indicated by the "All-Gather" arrow in the figure. This strategy distributes the encoding workload, allowing the system to handle significantly longer videos by leveraging more GPUs, while avoiding the risk of GPU memory overflow. The parallel encoding scheme ensures balanced utilization of the vision towers and enables scalable long-video processing that would otherwise be infeasible on a single device.
After the video embeddings are globally gathered, they are reused for downstream usage throughout the RL pipeline. As illustrated in Figure 7, the gathered embeddings are reused during multiple rollouts without recomputation. For instance, in each training step, we typically perform 8 to 16 rollouts. Without recycling, the same video would need to be re-encoded dozens of times per step, severely impacting training speed. By caching and reusing the gathered embeddings, MR-SP eliminates this redundancy and significantly accelerates training.
5.2 Stage 2 - Prefilling with Sequence Parallelism
For each rollout, both the reference and policy models require prefilling compute-intensively in RL for long videos. With the gathered embedding from Stage 1 reused, we parallelize the inference stage across devices using sequence parallelism. As illustrated in Figure 7, we globally gathered input embeddings are first padded to a uniform length (Padding Sequence) and then evenly partitioned across GPUs (Sharding to Local GPU). This allows each GPU to handle only a portion of the input sequence during prefilling. This parallelism is applied to both policy and reference model prefilling. Then, each GPU locally computes logits for its token slice, prefilling in parallel.
6 Experimental Results
6.1 Main Results
Table 1 shows the performance comparison on 9 video benchmarks [43, 44, 45, 46, 14, 47, 13, 48, 18]. LongVILA-R1-7B consistently outperforms LongVILA-7B across all benchmarks, with performance gaps that vary according to the complexity of the reasoning tasks. Table 3 presents the general performance of LongVILA-R1, comparing to existing advanced models [50, 57, 52, 58, 53, 59, 54, 60, 1, 2] under comparable model sizes on the Video-MME [18] benchmark. The LongVILA-R1-7B achieves the leading scores under different video lengths. LongVILA-R1-7B achieves leading scores across different video lengths, obtaining scores of 60.3 and 65.9 in the settings without subtitles and with subtitles, respectively. Table 2 compares the results of our LongVideo-Reason-eval benchmark. We also show some qualitative results in the appendix. LongVILA-R1-7B model achieves a strong performance with an average score of 67.9, significantly surpassing Video-R1-7B [2] and GPT-4o [19] and slightly outperforms Gemini-1.5-Pro [20]. While Gemini-1.5-Pro demonstrates strong performance in Temporal Reasoning, Goal and Purpose Reasoning, and Plot and Narrative Reasoning, the LongVILA-R1-7B reaches scores that are generally comparable. Notably, in the Spatial Reasoning category, LongVILA-R1-7B achieves a score of 70.0.
Figure 8: Scaling video frames improves long video reasoning performance against non-reasoning baselines.
6.2 Ablation Study
Scaling video frames.
The reasoning capability of LongVILA-R1 scales consistently with the number of input video frames. Specifically, Figure 8 illustrates the performance of LongVILA-1.5B (grey line) and LongVILA-1.5B-R1 (red line) on the long-video reasoning benchmark under varying frame inputs. With only 16 input frames, LongVILA-R1-1.5B is close to LongVILA-1.5B. However, as the number of frames increases to 512, LongVILA-R1-1.5B consistently outperforms and eventually achieves a score of 64.3. Notably, LongVILA-R1-1.5B demonstrates steady performance improvements throughout the scaling process. In contrast, LongVILA-1.5B hits a performance bottleneck with 256 frames, and got a degradation on 512 frames, due to its inability to effectively process such dense visual information. The enhanced reasoning capabilities of LongVILA-R1-1.5B allow it to effectively integrate and infer information from long videos.
Ablation on pipeline and datasets.
As shown in Table 4, we ablate the effectiveness of training stages and datasets, using LongVILA-1.5B as a starting point. The accuracies are evaluated on LongVideo-Reason-eval. ✗  means skipping this stage, ✓  means training this stage with our datasets, and O means training this stage with other datasets [12, 13, 14, 15, 16]. Our CoT-SFT dataset results in a better performance than that of other datasets. In addition, incorporating RL on top of the warm-up phase (CoT-SFT) yields additional improvements compared to using only SFT. We show that if we skip CoT-SFT and train our models directly with RL, the accuracy drops. If we apply Video-R1 datasets in both CoT-SFT and RL stages, the performance is inferior to using ours.
Training efficiency on MR-SP.
We conduct the training efficiency comparison for our MR-SP system, one A100 node, i.e., 8xA100 (80GB) GPUs. We measure the forward time for each training step. The results are obtained after 10 warming up iterations and averaged over 5 iterations to minimize variance. We use LongVILA-7B-R1 model with training batch size as 1 per GPU and rollout number as 5.
Figure 1 presents the training efficiency comparison across different numbers of frames. The figure plots the runtime per step (in seconds) for three settings: the plain RL system without MR-SP, MR-SP Stage 1 only, and the full MR-SP system (Stage 1 & 2). The baseline runtime increases steeply as the frame count grows. Using only Stage 1 of MR-SP significantly improves efficiency up to 512 frames but encounters GPU out-of-memory (OOM) issues beyond that point. In contrast, the full MR-SP system consistently reduces runtime, achieving up to a 2.1× speedup at 512 frames and scaling efficiently to 1024 frames without OOM, highlighting the benefit of combining sequence reuse and sequence parallelism for long video RL training.
Table 4: Ablations on training pipeline and datasets.
CoT-SFT O O
RL O
Accuracy 58.1 60.2 52.4 59.1 59.4 61.9
7 Conclusion
We present a comprehensive framework designed to fully scale VLMs for reasoning over long videos. LongVILA-R1 encompasses a meticulously constructed large-scale dataset, LongVideo-Reason, and a parallelized training framework, MR-SP. Leveraging our curated dataset of 52K long video question-reasoning-answer pairs, combined with other open-source video datasets, we adopt a two-stage training process that integrates CoT-SFT and RL. LongVILA-R1-7B demonstrates outstanding performance on mainstream video benchmarks, achieving 68.4% on VideoMME with subtitles, and attaining an average accuracy of 67.9% across four reasoning categories in the LongVideo-Reason testing set. LongVILA-R1 maintains consistent performance improvements, extending the feasible frame count from 16 to 512. Notably, our MR-SP leads to a speed up of 2.1$\times$ for long video RL training, supporting hour-level (3,600 frames / 256k tokens) RL training on a single node of 8 A100 GPUs. In addition, we release our training system to the public, which supports RL training across multiple modalities (video, text, and audio), various models (including the VILA and Qwen series), and even image and video generation models. In the appendix, we discuss limitations, border impacts, detailed data generation process, and provide more examples.
Appendix A Limitations
Despite our best efforts to create a high-quality long video reasoning dataset, the definition of reasoning still requires further refinement and a more comprehensive conclusion. Moreover, while our proposed efficient reasoning RL training framework is the first to support video training with hundreds to thousands of frames, real-world scenarios often involve videos with far greater frame counts. Thus, there remains significant potential for improvement in the extraction and learning of ultra-dense visual information.
Appendix B Border Impacts
The development of LongVILA-R1 represents a significant advancement in long video reasoning, enabling real world systems to perform sophisticated temporal and compositional understanding across diverse and extended visual contexts. This progress holds transformative potential across multiple domains, including embodied AI, robotics, autonomous systems, and AR/VR applications. By equipping vision-language models (VLMs) with the ability to process and reason over long-duration video data, LongVILA-R1 lays the foundation for AI systems capable of understanding event sequences, tracking object persistence and transformations, and inferring causal and physical relationships over extended frames. Long video reasoning technologies powered by LongVILA-R1 can significantly enhance embodied AI and robotics by enabling agents to sustain coherent, long-term understanding of their environment. Robots equipped with such capabilities would excel in performing complex, multi-stage tasks, adapting to dynamic contexts, and building richer world models for planning and decision-making. These advancements also promise to unlock new opportunities in education, healthcare, and entertainment. For instance, long video understanding could enable AI tutors to analyze and summarize extended instructional videos, or assist healthcare professionals in reviewing lengthy procedural recordings. Furthermore, such systems could enhance sports analytics, and other areas requiring nuanced temporal reasoning.
In conclusion, LongVILA-R1 demonstrates the potential of long video reasoning to drive progress across a wide range of applications, from robotics to immersive virtual environments. However, unlocking the full promise of this technology requires a steadfast commitment to ethical principles, privacy protection, and the broader goal of benefiting humanity. By addressing these challenges, the AI community can ensure that advancements in long video reasoning contribute positively to society while mitigating associated risks.
Appendix C More Demos in Different Reasoning Subsets
In this section, we selected examples of four types of reasoning tasks from the benchmark to evaluate and compare the reasoning processes and answers provided by Gemini-1.5-Pro, Video-R1-7B, and LongVILA-R1-7B. On Figure 12, a 20-minute StarCraft match is depicted, where the models analyze the players’ unit compositions, strategies, and play styles to predict the potential developments on the battlefield. While Gemini-1.5-Pro produced a correct prediction of the outcome, its reasoning process contained factual inaccuracies. In contrast, Video-R1-7B, influenced by the characteristics of its training data, tended to summarize answers based on options, neglecting critical video details and resulting in incorrect reasoning. LongVILA-R1-7B, however, is able to accurately analyze the players’ operational styles and specific moments marked in the video, leading to a comprehensive and accurate prediction of the match’s trajectory. On Figure 13, another example demonstrates the models’ abilities in narrative reasoning and visual information analysis. Gemini-1.5-Pro failed to correctly infer why the man appearing for the second time in the video is not the husband. In contrast, both Video-R1-7B and LongVILA-R1-7B successfully reasoned that the man’s habit of wearing a ring on his left hand is a key indicator, providing accurate answers.
Figure 9: Detailed data generation process for the LongVideo-Reason dataset, supplementing Figure 4.
Figure 10: Detailed comparisons in the football game example. The video is available at Link.
Figure 11: Detailed comparisons in the moving cup example. The video is available at Link.
Figure 12: LongVILA-R1 reasoning example in the "Goal and Purpose" category. The video is available at Link.
Figure 13: LongVILA-R1 reasoning example in the "Plot and Narrative" category. The video is available at Link.
Figure 14: LongVILA-R1 reasoning example in the "Spatial" category. The video is available at Link.
Figure 15: LongVILA-R1 reasoning example in the "Temporal" category. The video is available at Link.
Figure 16: LongVILA-R1 reasoning example in the taboo game. The video is available at Link.
Figure 14 illustrates the models’ spatial perception and reasoning abilities as the camera moves through a room. Gemini-1.5-Pro effectively identified the key information within the video and provided the correct answer through straightforward reasoning. In contrast, Video-R1-7B experienced significant localization errors during the reasoning process, leading to a critical issue for reasoning models: a mismatch between the reasoning analysis and the final answer. LongVILA-R1-7B demonstrated superior performance by leveraging dense frame analysis to accurately infer the spatial relationships between rooms and furniture across different levels, ultimately delivering a coherent reasoning process and the correct answer. On Figure 15, the focus shifts to temporal analysis in a Lego video featuring diverse events and interactions. All three models successfully reasoned through the sequence of events and provided correct answers, showcasing their proficiency in temporal reasoning tasks. As a supplement to Figure 2, Figure 10 provides a more comprehensive comparison of two examples: "2022 FIFA Argentina vs. Netherlands" and "Moving the Cup and Guessing Where the Ball Is." In the football match example, while Gemini-1.5-Pro produced the correct answer, its output contained hallucinatory content influenced by biases in its pre-learned knowledge. Video-R1 not only failed to provide accurate video analysis reasoning but also made incorrect predictions. In contrast, LongVILA-R1 successfully analyzed the players’ performance and emotions during the match, integrating these factors through its robust reasoning capabilities to make accurate predictions about the outcome. For the more challenging task of tracking the ball, Gemini-1.5-Pro’s reasoning is inconsistent with the spatial content throughout, while Video-R1 failed to deduce the ball’s final position accurately. Remarkably, LongVILA-R1 precisely analyzed the spatial transformations following the movement of the box, demonstrating superior interpretative and reasoning abilities.

### v2

@@
 Abstract: We introduce a full-stack framework that scales up reasoning in vision-language models (VLMs) to long videos, leveraging reinforcement learning.
-We address the unique challenges of long video reasoning by integrating three critical components: (1) a large-scale dataset, LongVideo-Reason, comprising 52K long video QA pairs with high-quality reasoning annotations across diverse domains such as sports, games, and vlogs; (2) a two-stage training pipeline that extends VLMs with chain-of-thought supervised fine-tuning (CoT-SFT) and reinforcement learning (RL); and (3) a training infrastructure for long video RL, named Multi-modal Reinforcement Sequence Parallelism (MR-SP), which incorporates sequence parallelism and a vLLM-based engine tailored for long video, using cached video embeddings for efficient rollout and prefilling.
-In experiments, LongVILA-R1-7B achieves strong performance on long video QA benchmarks such as VideoMME.
-It also outperforms Video-R1-7B and even matches Gemini-1.5-Pro across temporal reasoning, goal and purpose reasoning, spatial reasoning, and plot reasoning on our LongVideo-Reason-eval benchmark.
+We address the unique challenges of long video reasoning by integrating three critical components: (1) a large-scale dataset, LongVideo-Reason, comprising 104K long video QA pairs with high-quality reasoning annotations across diverse domains such as sports, games, and vlogs; (2) a two-stage training pipeline that extends VLMs with chain-of-thought supervised fine-tuning (CoT-SFT) and reinforcement learning (RL); and (3) a training infrastructure for long video RL, named Multi-modal Reinforcement Sequence Parallelism (MR-SP), which incorporates sequence parallelism and a vLLM-based engine tailored for long video, using cached video embeddings for efficient rollout and prefilling.
+In our experiments, LongVILA-R1-7B achieves strong performance on video benchmarks, reaching 65.0% and 70.7% accuracy on VideoMME without and with subtitles, respectively, and consistently outperforming LongVILA-R1 across multiple benchmarks.
+Moreover, LongVILA-R1 shows steady performance improvements as the number of input video frames increases.
 Notably, our MR-SP system achieves up to 2.1$\times$ speedup on long video RL training.
-LongVILA-R1 demonstrates consistent performance gains as the number of input video frames scales.
-LongVILA-R1 marks a firm step towards long video reasoning in VLMs.
 In addition, we release our training system for public availability that supports RL training on various modalities (video, text, and audio), various models (VILA and Qwen series), and even image and video generation models.
 On a single A100 node (8 GPUs), it supports RL training on hour-long videos (e.g., 3,600 frames / around 256k tokens).
-Code and models are available at https://github.com/NVlabs/Long-RL.
+Code and models are available at https://github.com/NVlabs/Long-RL and https://huggingface.co/Efficient-Large-Model/LongVILA-R1-7B.
 1 Introduction
@@
 Firstly, we strategically construct a high-quality dataset with CoT annotations for long video reasoning, named LongVideo-Reason.
-Leveraging a powerful VLM (NVILA-8B) [11] and a leading open-source reasoning LLM, we develop a dataset comprising 52K high-quality Question-Reasoning-Answer pairs for long videos.
-We use 18K high-quality samples for Long-CoT-SFT to initialize the model’s reasoning and instruction-following abilities, and 33K samples with an additional 110K video data [2] for reinforcement learning.
+Leveraging a powerful VLM (NVILA-8B) [11] and a leading open-source reasoning LLM, we develop a dataset comprising 104K high-quality Question-Reasoning-Answer pairs for long videos.
+We use 36K high-quality samples for Long-CoT-SFT to initialize the model’s reasoning and instruction-following abilities, and 68K samples with an additional 102K video data [12, 13, 14, 15, 16] for reinforcement learning.
 This two-stage training combines high-quality reasoning annotations with reinforcement learning, enabling LongVILA-R1 to achieve superior and generalized video reasoning.
@@
 Figure 3: Data Distribution of LongVideo-Reason and total data in the LongVILA-R1 training framework.
-LongVideo-Reason comprises a total of 18,077 videos, from which 52K QAs with reasoning annotations.
-Additionally, we also include additional 110K QAs from existing works [12, 13, 14, 15, 16] for RL training.
+LongVideo-Reason comprises a total of 18K videos, from which 104K QAs with reasoning annotations.
+Additionally, we also include additional 102K QAs from existing works [12, 13, 14, 15, 16] for RL training.
 Secondly, we propose a training framework for VLMs to advance long video reasoning.
@@
 As shown in Figure 1, our MR-SP system achieves up to 2.1$\times$ speedup on 512-frame video RL training on 7B models and enables longer training frames without out-of-memory (OOM).
-LongVILA-R1-7B shows strong performance on general video benchmark, e.g., 68.4% on VideoMME [18](with subtitle).
-On our LongVideo-Reason-eval benchmark, LongVILA-R1-7B achieves an average accuracy of 67.9% across four reasoning categories, surpassing open-source models including Video-R1-7B [2] and proprietary models such as GPT-4o [19] by a large margin, and matches the performance of Gemini-1.5-Pro [20].
-In addition, LongVILA-R1 exhibits progressive performance improvements as the number of input frames increased.
+In our experiments, LongVILA-R1-7B demonstrates strong performance on video benchmarks, achieving 65.0% and 70.7% accuracy on VideoMME without and with subtitles, respectively.
+It consistently outperforms LongVILA-R1 across a range of benchmarks, including ActivityNet-QA, LongVideoBench, PerceptionTest, NExT-QA, VNBench, and VideoMME.
+Additionally, LongVILA-R1 exhibits steady performance gains as the number of input video frames increases from 16 to 512 frames.
+On our LongVideo-Reason-eval benchmark, LongVILA-R1-7B achieves an average accuracy of 72.0% across four reasoning categories—surpassing open-source models like Video-R1-7B[2], as well as proprietary models such as Gemini-1.5-Pro[18].
+Figure 4: Data generation process for the LongVideo-Reason dataset.
+This process begins with segmenting videos into 10-second clips and generating captions for each clip using NVILA-8B.
+Then based on the captions of all clips in a video, we generate question-answer pairs that involve reasoning across the content of the whole video, along with the reasoning annotations using a leading open-source reasoning LLM.
+Reasoning questions are categorized into Temporal, Goal and Purpose, Spatial, and Plot and Narrative.
+Finally, the reasoning annotations are reformatted for conciseness and alignment with video details.
+We present a more detailed figure of data generation process in Figure 9 in the appendix.
 2 Related Work
@@
 In this work, we propose MR-SP that provides up to 2.1$\times$ speedup for long video RL training.
-Figure 4: Data generation process for the LongVideo-Reason dataset.
-This process begins with segmenting videos into 10-second clips and generating captions for each clip using NVILA-8B.
-Then based on the captions of all clips in a video, we generate question-answer pairs that involve reasoning across the content of the whole video, along with the reasoning annotations using a leading open-source reasoning LLM.
-Reasoning questions are categorized into Temporal, Goal and Purpose, Spatial, and Plot and Narrative.
-Finally, the reasoning annotations are reformatted for conciseness and alignment with video details.
-We present a more detailed figure of data generation process in Figure 9 in the appendix.
 3 LongVideo-Reason Data Construction
@@
 Existing publicly available long-video datasets lack high-quality reasoning annotations.
-In this section, we detail the data construction process (Figure 4) of a 52K long-video QAs with reasons.
+In this section, we detail the data construction process (Figure 4) of a 104K long-video QAs with reasons.
 3.1 Overview of Data Curation
 We first curate 18K long videos from the Shot2Story dataset [40] (Figure 3, left).
-We then apply a high-quality automated annotation pipeline for CoT as detailed in Section 3.2, and end up with a total of 52K Question-Reasoning-Answer pairs where each sample, based on the type of question it is reasoning about, can be categorized into Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning (Figure 3, middle).
+We then apply a high-quality automated annotation pipeline for CoT as detailed in Section 3.2, and end up with a total of 104K Question-Reasoning-Answer pairs where each sample, based on the type of question it is reasoning about, can be categorized into Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning (Figure 3, middle).
 This dataset is designed to support various types of long-video reasoning tasks comprehensively.
@@
 Specifically, we employ a test-scaling method where LongVILA [4] performs inference 10 times on the original datasets.
-Questions consistently answered correctly or incorrectly are labeled as "easy" or "hard" while those inducing diverse predictions are labeled as "medium".
-We use both the "easy" and "hard" samples for Stage-1 COT-SFT and use the "medium" samples for Stage-2 RL training.
+Questions consistently answered correctly or incorrectly are labeled as easy or hard while those inducing diverse predictions are labeled as medium.
+We filter out too easy and too hard questions, and use remaining questions for training.
 The reason is that GRPO expects different rollouts of each sample to be diverse in order to have meaningful advantages, and the gradient vanishes if all the rollouts predict correct or incorrect answers.
-The COT-SFT subset (18K) features high-quality CoT reasoning processes formatted in a standard <think></think><answer></answer> structure, providing abundant resources for warm-up training during Stage 1 of the model’s reasoning capabilities.
-Meanwhile, the RL subset contains 33K challenging long-video Q&A, which is leveraged in Stage 2 for scaling reasoning through reinforcement learning.
-To further enhance RL scaling, we incorporate an additional 110K high-quality open-source videos (Figure 3, right) from other datasets [12, 13, 14, 15, 16].
+The COT-SFT subset (36K) features high-quality CoT reasoning processes formatted in a standard <think></think><answer></answer> structure, providing abundant resources for warm-up training during Stage 1 of the model’s reasoning capabilities.
+Meanwhile, the RL subset contains 68K challenging long-video Q&A, which is leveraged in Stage 2 for scaling reasoning through reinforcement learning.
+To further enhance RL scaling, we incorporate an additional 102K high-quality open-source videos (Figure 3, right) from other datasets [12, 13, 14, 15, 16].
 This combination improves the model’s generalization.
@@
 We also manually curated 1,000 high-quality complex reasoning questions across four reasoning categories to serve as a new benchmark (LongVideo-Reason-eval) for evaluating VLMs in reasoning abilities.
-This entire data procedure consumes about 40,000 GPU hours on H100s.
+Note that we also include open-ended questions.
+Among the total 104K QA pairs, approximately half are multiple-choice and the other half are open-ended.
+This entire data procedure consumes about 80,000 GPU hours on H100s.
 4 LongVILA Training Pipeline
-As shown in Figure 5, there are two extended training stages in LongVILA-R1, i.e., (1) warm-up for long video reasoning, utilizing 18K data with high-quality CoT for SFT on the MM-SP system [4]; (2) reinforcement learning with dense frames from long videos.
+As shown in Figure 5, there are two extended training stages in LongVILA-R1, i.e., (1) warm-up for long video reasoning, utilizing 36K data with high-quality CoT for SFT on the MM-SP system [4]; (2) reinforcement learning with dense frames from long videos.
 4.1 Long Video CoT Supervised Fine-Tuning
-Utilizing 52K high-quality question-reasoning-answer pairs, we apply the data filtering method described in Section 3.1 to select 18K examples for long CoT-SFT, serving as a warm-up phase for subsequent RL.
+Utilizing 104K high-quality question-reasoning-answer pairs, we apply the data filtering method described in Section 3.1 to select 36K examples for long CoT-SFT, serving as a warm-up phase for subsequent RL.
 This stage equips the model with fundamental reasoning abilities and instruction-following skills for long video scenarios.
@@
 To address this, we develop the MR-SP framework (Section 5), which efficiently scales reinforcement learning for long-context video reasoning.
-Considering the sensitivity of GRPO to sampling during training [41], we use the 33K filtered data for reinforcement learning as described in Section 3.1.
-Additionally, an extra 110K samples from Video-R1 [2] are incorporated to scale up the RL.
+Considering the sensitivity of GRPO to sampling during training [41], we use the 68K filtered data for reinforcement learning as described in Section 3.1.
+Additionally, an extra 102K samples from other datasets [12, 13, 14, 15, 16] are incorporated to scale up the RL.
 This approach aims to guide the model in freely exploring and developing more effective and generalized reasoning strategies.
@@
 MR-SP leverages sequence parallelism in both rollout and pre-filling stages, enabling long videos in RL, with reduced overhead.
-Table 1: Performance on 9 video benchmarks [43, 44, 45, 46, 14, 47, 13, 48, 18].
-LongVILA-R1-7B outperforms LongVILA-7B across all benchmarks, with varying margins that reflect different levels of reasoning difficulty.
-Model
-ActNet-QA
-EgoSchema
-EventBench
-LVideoBench
-PercepTest
-MVBench
-NExT-QA
-VNBench
-VideoMME
-test test val val val test mc val w/o sub. w/ sub.
-GPT-4V [49] 57.0 - 32.6 61.3 - 43.5 - - 59.9 63.3
-GPT-4o [19] - - 53.3 66.7 - - - 64.4 71.9 77.2
-Gemini-1.5-Pro [20] 57.5 72.2 43.2 64.0 - - - 66.7 75.0 81.3
-Video-LLaVA-7B [50] 45.3 38.4 5.9 37.6 - 43.5 - 12.4 39.9 41.6
-Flash-VStream-7B [51] 51.9 - - - - - 61.6 - - -
-ShareGPT4Video-8B [52] 50.8 - - 41.8 - 51.2 - - 39.9 43.6
-VideoLLaMA2-7B [53] 50.2 51.7 6.9 - 51.4 54.6 - 4.5 47.9 50.3
-VideoLLaMA2.1-7B [53] 53.0 53.1 - - 54.9 57.3 - - 54.9 56.4
-Kangaroo-8B [54] - 62.7 - 54.8 - 61.1 - - 56.0 57.6
-PLLaVA-7B [55] 56.3 - 28.2 39.2 - 46.6 - - - -
-LLaVA-OV-7B [56] 56.7 60.1 - 56.4 57.1 56.7 79.4 51.8 58.2 61.5
-LongVILA-7B [4] 59.5 67.7 58.0 57.1 58.1 67.1 80.7 63.0 60.1 65.1
-LongVILA-R1-7B 59.6 68.1 58.2 57.6 59.3 67.6 80.8 66.8 62.4 68.4
-Table 2: Performance on LongVideo-Reason-eval. LongVILA-R1-7B achieves the highest overall score.
-     Model      Temporal      Goal      Plot      Spatial      Overall
-     GPT-4o [19]      64.4      62.4      60.4      55.6      60.7
-     Gemini-1.5-Pro [20]      73.2      66.4      66.4      63.2      67.3
-     Video-R1-7B [2]      66.4      62.8      63.6      58.0      62.7
-     LongVILA-R1-7B      71.6      66.4      63.6      70.0      67.9
-Table 3: Performance comparison on VideoMME [18] benchmark in details.
-Model w/o subtitle w subtitle
-Overall Short Medium Long Overall Short Medium Long
-Video-LLaVA-7B [50] 39.9 45.3 38.0 36.2 41.6 46.1 40.7 38.1
-SliME-8B [57] 45.3 53.3 55.4 39.8 47.2 55.4 44.4 41.7
-ShareGPT4Video-8B [52] 39.9 48.3 36.3 35.0 43.6 53.6 39.3 37.9
-VideoChat2-7B [58] 39.5 48.3 37.0 33.2 43.8 52.8 39.4 39.2
-VideoLLaMA2-7B [53] 47.9 56.0 45.4 42.1 50.3 59.4 47.6 43.8
-Chat-Univi-v1.5-7B [59] 40.6 45.7 40.3 35.8 45.9 51.2 44.6 41.8
-Kangaroo-8B [54] 56.0 66.1 55.3 46.7 57.6 68.0 55.4 49.3
-ShareGemini-7B [60] 43.2 49.1 41.3 39.1 47.9 49.1 47.3 43.4
-LongVA-7B [1] 52.6 61.1 50.4 46.2 54.3 61.1 53.6 47.6
-Video-R1-7B [2] 59.3 - - - - - - -
-LongVILA-7B [4] 60.1 69.0 58.3 53.0 65.1 72.9 64.9 57.4
-LongVILA-R1-7B 62.4 71.9 62.0 53.3 68.4 76.2 69.8 59.1
 5.1 Stage 1 - Rollout with Paralleled Encoding
@@
 Then, each GPU locally computes logits for its token slice, prefilling in parallel.
+Table 1: Performance on 6 video benchmarks [43, 44, 14, 13, 45, 20].
+LongVILA-R1-7B outperforms LongVILA-7B across all benchmarks, with varying margins that reflect different levels of reasoning difficulty.
+Model ActivityNet-QA LongVideoBench PerceptionTest NExT-QA VNBench VideoMME
+test val val mc val w/o sub. w/ sub.
+GPT-4V 57.0 61.3 - - - 59.9 63.3
+GPT-4o - 66.7 - - 64.4 71.9 77.2
+Gemini-1.5-Pro 57.5 64.0 - - 66.7 75.0 81.3
+Video-LLaVA-7B 45.3 37.6 - - 12.4 39.9 41.6
+Flash-VStream-7B 51.9 - - 61.6 - - -
+ShareGPT4Video-8B 50.8 41.8 - - - 39.9 43.6
+VideoLLaMA2-7B 50.2 - 51.4 - 4.5 47.9 50.3
+VideoLLaMA2.1-7B 53.0 - 54.9 - - 54.9 56.4
+Kangaroo-8B - 54.8 - - - 56.0 57.6
+PLLaVA-7B 56.3 39.2 - - - - -
+LLaVA-OV-7B 56.7 56.4 57.1 79.4 51.8 58.2 61.5
+LongVILA-7B 59.5 57.1 58.1 80.7 63.0 60.1 65.1
+LongVILA-R1-7B 64.8 58.0 68.9 81.5 75.5 65.0 70.7
+Table 2: Performance on LongVideo-Reason-eval set. LongVILA-R1-7B achieves the highest overall score.
+     Model      Temporal      Goal      Plot      Spatial      Overall
+     Video-R1-7B [2]      61.4      85.0      62.0      58.5      68.1
+     Gemini-1.5-Pro [18]      65.4      81.9      67.8      53.3      69.3
+     LongVILA-R1-7B      68.1      85.7      70.6      53.3      72.0
+Table 3: Performance comparison on VideoMME [20] benchmark in details.
+Model w/o subtitle w subtitle
+Overall Short Medium Long Overall Short Medium Long
+Video-LLaVA-7B [46] 39.9 45.3 38.0 36.2 41.6 46.1 40.7 38.1
+SliME-8B [47] 45.3 53.3 55.4 39.8 47.2 55.4 44.4 41.7
+ShareGPT4Video-8B [48] 39.9 48.3 36.3 35.0 43.6 53.6 39.3 37.9
+VideoChat2-7B [49] 39.5 48.3 37.0 33.2 43.8 52.8 39.4 39.2
+VideoLLaMA2-7B [50] 47.9 56.0 45.4 42.1 50.3 59.4 47.6 43.8
+Chat-Univi-v1.5-7B [51] 40.6 45.7 40.3 35.8 45.9 51.2 44.6 41.8
+Kangaroo-8B [52] 56.0 66.1 55.3 46.7 57.6 68.0 55.4 49.3
+ShareGemini-7B [53] 43.2 49.1 41.3 39.1 47.9 49.1 47.3 43.4
+LongVA-7B [1] 52.6 61.1 50.4 46.2 54.3 61.1 53.6 47.6
+Video-R1-7B [2] 61.4 - - - - - - -
+LongVILA-7B [4] 60.1 69.0 58.3 53.0 65.1 72.9 64.9 57.4
+LongVILA-R1-7B 65.0 76.6 63.6 54.9 70.7 79.0 69.1 64.0
 6 Experimental Results
 6.1 Main Results
-Table 1 shows the performance comparison on 9 video benchmarks [43, 44, 45, 46, 14, 47, 13, 48, 18].
+Table 1 shows the performance comparison on 6 video benchmarks [43, 44, 14, 13, 45, 20].
 LongVILA-R1-7B consistently outperforms LongVILA-7B across all benchmarks, with performance gaps that vary according to the complexity of the reasoning tasks.
 Table 3 presents the general performance of LongVILA-R1, comparing to existing advanced models [46, 47, 48, 49, 50, 51, 52, 53, 1, 2] under comparable model sizes on the Video-MME [20] benchmark.
-The LongVILA-R1-7B achieves the leading scores under different video lengths.
-LongVILA-R1-7B achieves leading scores across different video lengths, obtaining scores of 60.3 and 65.9 in the settings without subtitles and with subtitles, respectively.
+The LongVILA-R1-7B is tested using 512 video frames as inputs.
+LongVILA-R1-7B achieves leading scores across different video lengths, obtaining scores of 65.0% and 70.7% in the settings without subtitles and with subtitles, respectively.
 Table 2 compares the results of our LongVideo-Reason-eval benchmark.
 We also show some qualitative results in the appendix.
-LongVILA-R1-7B model achieves a strong performance with an average score of 67.9, significantly surpassing Video-R1-7B [2] and GPT-4o [19] and slightly outperforms Gemini-1.5-Pro [20].
-While Gemini-1.5-Pro demonstrates strong performance in Temporal Reasoning, Goal and Purpose Reasoning, and Plot and Narrative Reasoning, the LongVILA-R1-7B reaches scores that are generally comparable.
-Notably, in the Spatial Reasoning category, LongVILA-R1-7B achieves a score of 70.0.
+LongVILA-R1-7B model achieves a strong performance with an average score of 72.0%, significantly surpassing Video-R1-7B [2] and slightly outperforms Gemini-1.5-Pro [18].
 Figure 8: Scaling video frames improves long video reasoning performance against non-reasoning baselines.
@@
 With only 16 input frames, LongVILA-R1-1.5B is close to LongVILA-1.5B.
-However, as the number of frames increases to 512, LongVILA-R1-1.5B consistently outperforms and eventually achieves a score of 64.3.
+However, as the number of frames increases to 512, LongVILA-R1-1.5B consistently outperforms and eventually achieves a score of 64.3%.
 Notably, LongVILA-R1-1.5B demonstrates steady performance improvements throughout the scaling process.
@@
 Leveraging our curated dataset of 52K long video question-reasoning-answer pairs, combined with other open-source video datasets, we adopt a two-stage training process that integrates CoT-SFT and RL.
-LongVILA-R1-7B demonstrates outstanding performance on mainstream video benchmarks, achieving 68.4% on VideoMME with subtitles, and attaining an average accuracy of 67.9% across four reasoning categories in the LongVideo-Reason testing set.
+LongVILA-R1-7B demonstrates outstanding performance on mainstream video benchmarks, achieving 65.0% and 70.7% on VideoMME without and with subtitles.
 LongVILA-R1 maintains consistent performance improvements, extending the feasible frame count from 16 to 512.
@@
 In addition, we release our training system to the public, which supports RL training across multiple modalities (video, text, and audio), various models (including the VILA and Qwen series), and even image and video generation models.
-In the appendix, we discuss limitations, border impacts, detailed data generation process, and provide more examples.
-Appendix A Limitations
-Despite our best efforts to create a high-quality long video reasoning dataset, the definition of reasoning still requires further refinement and a more comprehensive conclusion.
-Moreover, while our proposed efficient reasoning RL training framework is the first to support video training with hundreds to thousands of frames, real-world scenarios often involve videos with far greater frame counts.
-Thus, there remains significant potential for improvement in the extraction and learning of ultra-dense visual information.
-Appendix B Border Impacts
+In the appendix, we discuss RL training curves, border impacts, detailed data generation process, and provide more examples.
+Limitations
+While our RL training system is capable of handling long video frames efficiently (3,600 frames on a single node of A100 GPUs), scaling to significantly longer sequences, more modalities (like include audio for omni VLMs) or large batch sizes would require distributed training across multiple GPU machines.
+This requires more GPU machines, making it less feasible to run on limited hardware resources.
+Appendix A Border Impacts
 The development of LongVILA-R1 represents a significant advancement in long video reasoning, enabling real world systems to perform sophisticated temporal and compositional understanding across diverse and extended visual contexts.
@@
 By addressing these challenges, the AI community can ensure that advancements in long video reasoning contribute positively to society while mitigating associated risks.
-Appendix C More Demos in Different Reasoning Subsets
+Appendix B More Demos in Different Reasoning Subsets
 In this section, we selected examples of four types of reasoning tasks from the benchmark to evaluate and compare the reasoning processes and answers provided by Gemini-1.5-Pro, Video-R1-7B, and LongVILA-R1-7B.
@@
 Figure 9: Detailed data generation process for the LongVideo-Reason dataset, supplementing Figure 4.
+Figure 10: Reward curve of our LongVILA-R1-7B. During training, LongVILA-R1-7B exhibits stable and consistent reward improvements across overall, format, and accuracy dimensions, converging steadily.
 Figure 11: Detailed comparisons in the football game example. The video is available at Link.

### v3

@@
 We address the unique challenges of long video reasoning by integrating three critical components: (1) a large-scale dataset, LongVideo-Reason, comprising 104K long video QA pairs with high-quality reasoning annotations across diverse domains such as sports, games, and vlogs; (2) a two-stage training pipeline that extends VLMs with chain-of-thought supervised fine-tuning (CoT-SFT) and reinforcement learning (RL); and (3) a training infrastructure for long video RL, named Multi-modal Reinforcement Sequence Parallelism (MR-SP), which incorporates sequence parallelism and a vLLM-based engine tailored for long video, using cached video embeddings for efficient rollout and prefilling.
-In our experiments, LongVILA-R1-7B achieves strong performance on video benchmarks, reaching 65.0% and 70.7% accuracy on VideoMME without and with subtitles, respectively, and consistently outperforming LongVILA-R1 across multiple benchmarks.
-Moreover, LongVILA-R1 shows steady performance improvements as the number of input video frames increases.
+In our experiments, LongVILA-R1-7B achieves strong performance on video benchmarks, reaching 65.1% and 71.1% accuracy on VideoMME without and with subtitles, respectively, and consistently outperforming LongVILA-7B across multiple benchmarks.
+Moreover, LongVILA-R1-7B supports processing up to 8,192 video frames per video, and configurable FPS settings.
 Notably, our MR-SP system achieves up to 2.1$\times$ speedup on long video RL training.
 In addition, we release our training system for public availability that supports RL training on various modalities (video, text, and audio), various models (VILA and Qwen series), and even image and video generation models.
-On a single A100 node (8 GPUs), it supports RL training on hour-long videos (e.g., 3,600 frames / around 256k tokens).
+On a single A100 node (8 GPUs), it supports RL training on hour-long videos (e.g., 3,600 frames).
 Code and models are available at https://github.com/NVlabs/Long-RL and https://huggingface.co/Efficient-Large-Model/LongVILA-R1-7B.
@@
 Notably, the spatial tracking video involves a relatively complex dynamic moving, for which the model fails to achieve accurate reasoning until the number of input video frames increases to 128.
-Despite the clear importance of reasoning in long video understanding, enabling such capabilities in long-video VLMs poses significant challenges [2, 3, 4, 5, 6].
-First, the collection of high-quality long video reasoning datasets is inherently difficult.
+Despite the clear importance of reasoning in long video understanding, enabling such capabilities in long video VLMs poses significant challenges [2, 3, 4, 5, 6].
+First, the collection of high quality long video reasoning datasets is inherently difficult.
 Unlike domains such as math or code reasoning, where structured supervision and benchmarks are readily available [7, 8], long video reasoning requires annotating complex temporal dynamics, goals, spatial relations, and narrative elements—often across minutes or hours of footage [9].
@@
 These challenges jointly hinder the development of effective long video VLMs with strong reasoning capabilities.
-In this work, we introduce LongVILA-R1, a comprehensive framework exploring the reasoning capabilities for long-video understandings.
-Firstly, we strategically construct a high-quality dataset with CoT annotations for long video reasoning, named LongVideo-Reason.
-Leveraging a powerful VLM (NVILA-8B) [11] and a leading open-source reasoning LLM, we develop a dataset comprising 104K high-quality Question-Reasoning-Answer pairs for long videos.
-We use 36K high-quality samples for Long-CoT-SFT to initialize the model’s reasoning and instruction-following abilities, and 68K samples with an additional 102K video data [12, 13, 14, 15, 16] for reinforcement learning.
-This two-stage training combines high-quality reasoning annotations with reinforcement learning, enabling LongVILA-R1 to achieve superior and generalized video reasoning.
-We also manually curate a balanced set of 1K long-video samples to build a new benchmark, LongVideo-Reason-eval, that evaluates performance from four perspectives: Temporal, Goal and Purpose, Spatial, and Plot and Narrative, for a comprehensive assessment.
+In this work, we introduce LongVILA-R1, a comprehensive framework exploring the reasoning capabilities for long video understandings.
+Firstly, we strategically construct a high quality dataset with CoT annotations for long video reasoning, named LongVideo-Reason.
+Leveraging a powerful VLM (NVILA-8B) [11] and a leading open-source reasoning LLM, we develop a dataset comprising 104K high quality Question-Reasoning-Answer pairs for long videos.
+We use 36K high quality samples for Long-CoT-SFT to initialize the model’s reasoning and instruction-following abilities, and 68K samples with an additional 102K video data [12, 13, 14, 15, 16] for reinforcement learning.
+This two-stage training combines high quality reasoning annotations with reinforcement learning, enabling LongVILA-R1 to achieve superior and generalized video reasoning.
+We also manually curate a balanced set of 1K long video samples to build a new benchmark, LongVideo-Reason-eval, that evaluates performance from four perspectives: Temporal, Goal and Purpose, Spatial, and Plot and Narrative, for a comprehensive assessment.
 Figure 3: Data Distribution of LongVideo-Reason and total data in the LongVILA-R1 training framework.
-LongVideo-Reason comprises a total of 18K videos, from which 104K QAs with reasoning annotations.
+LongVideo-Reason comprises a total of 18K videos and 104K QAs with reasoning annotations.
 Additionally, we also include additional 102K QAs from existing works [12, 13, 14, 15, 16] for RL training.
@@
 As shown in Figure 1, our MR-SP system achieves up to 2.1$\times$ speedup on 512-frame video RL training on 7B models and enables longer training frames without out-of-memory (OOM).
-In our experiments, LongVILA-R1-7B demonstrates strong performance on video benchmarks, achieving 65.0% and 70.7% accuracy on VideoMME without and with subtitles, respectively.
-It consistently outperforms LongVILA-R1 across a range of benchmarks, including ActivityNet-QA, LongVideoBench, PerceptionTest, NExT-QA, VNBench, and VideoMME.
-Additionally, LongVILA-R1 exhibits steady performance gains as the number of input video frames increases from 16 to 512 frames.
-On our LongVideo-Reason-eval benchmark, LongVILA-R1-7B achieves an average accuracy of 72.0% across four reasoning categories—surpassing open-source models like Video-R1-7B[2], as well as proprietary models such as Gemini-1.5-Pro[18].
+In our experiments, LongVILA-R1-7B demonstrates strong performance on video benchmarks, achieving 65.1% and 71.1% accuracy on VideoMME without and with subtitles, respectively.
+It consistently outperforms LongVILA-7B across a range of benchmarks, including ActivityNet-QA, LongVideoBench, PerceptionTest, NExT-QA, VNBench, and VideoMME.
+Additionally, LongVILA-R1 supports processing up to 8,192 video frames per video, and configurable FPS settings.
+On our LongVideo-Reason-eval benchmark, LongVILA-R1-7B achieves an average accuracy of 72.0%, surpassing open-source models like Video-R1-7B[2], as well as proprietary models such as Gemini-1.5-Pro[18].
 Figure 4: Data generation process for the LongVideo-Reason dataset.
@@
 3 LongVideo-Reason Data Construction
-High-quality annotated datasets for long videos are critical for VLMs.
-Existing publicly available long-video datasets lack high-quality reasoning annotations.
-In this section, we detail the data construction process (Figure 4) of a 104K long-video QAs with reasons.
 3.1 Overview of Data Curation
 We first curate 18K long videos from the Shot2Story dataset [40] (Figure 3, left).
-We then apply a high-quality automated annotation pipeline for CoT as detailed in Section 3.2, and end up with a total of 104K Question-Reasoning-Answer pairs where each sample, based on the type of question it is reasoning about, can be categorized into Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning (Figure 3, middle).
+We further incorporated 2k additional 4K-resolution videos spanning scenarios such as autonomous driving, video games, household robotics, and wildlife.
+We then apply a high quality automated annotation pipeline for CoT as detailed in Section 3.2, and end up with a total of 104K Question-Reasoning-Answer pairs where each sample, based on the type of question it is reasoning about, can be categorized into Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning (Figure 3, middle).
 This dataset is designed to support various types of long-video reasoning tasks comprehensively.
@@
 The reason is that GRPO expects different rollouts of each sample to be diverse in order to have meaningful advantages, and the gradient vanishes if all the rollouts predict correct or incorrect answers.
-The COT-SFT subset (36K) features high-quality CoT reasoning processes formatted in a standard <think></think><answer></answer> structure, providing abundant resources for warm-up training during Stage 1 of the model’s reasoning capabilities.
+The COT-SFT subset (36K) features high quality CoT reasoning processes formatted in a standard <think></think><answer></answer> structure, providing abundant resources for warm-up training during Stage 1 of the model’s reasoning capabilities.
 Meanwhile, the RL subset contains 68K challenging long-video Q&A, which is leveraged in Stage 2 for scaling reasoning through reinforcement learning.
-To further enhance RL scaling, we incorporate an additional 102K high-quality open-source videos (Figure 3, right) from other datasets [12, 13, 14, 15, 16].
+To further enhance RL scaling, we incorporate an additional 102K high quality open-source videos (Figure 3, right) from other datasets [12, 13, 14, 15, 16].
 This combination improves the model’s generalization.
@@
 3.2 Long-Video Reasoning Generation
-We introduce an automated annotation pipeline (Figure 4) that generates high-quality Question-Reasoning-Answer pairs from long videos.
+We introduce an automated annotation pipeline (Figure 4) that generates high quality Question-Reasoning-Answer pairs from long videos.
 This process begins by segmenting videos into short clips (~10 seconds each), each of which is annotated using the NVILA-8B [11] model to provide descriptive captions.
-Leveraging the breakthrough in text-based reasoning, we then deploy a leading open-source reasoning LLM, provide the captions of all the clips in each video, and prompt it to generate diverse types of Question-Reasoning-Answer pairs that involve reasoning over the content across the whole video.
+For the spatial reasoning category, We employed VILA-HD [24] to generate object bounding boxes in video frames and constructed spatial reasoning QAs based on these boxes and corresponding captions.
+Leveraging the breakthrough in text-based reasoning, we then deploy a leading open-source reasoning LLM, DeepSeek-R1-671B [36], provide the captions of all the clips in each video, and prompt it to generate diverse types of Question-Reasoning-Answer pairs that involve reasoning over the content across the whole video.
 Specifically, we design four types of prompts to encourage the LLM to generate a Question-Reasoning-Answer pair that focuses on one of the four types of reasoning: Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning.
@@
 Finally, an LLM is then used to refine and streamline the reasoning steps.
-We also manually curated 1,000 high-quality complex reasoning questions across four reasoning categories to serve as a new benchmark (LongVideo-Reason-eval) for evaluating VLMs in reasoning abilities.
+We also manually curated 1,000 high quality complex reasoning questions across four reasoning categories to serve as a new benchmark (LongVideo-Reason-eval) for evaluating VLMs in reasoning abilities.
 Note that we also include open-ended questions.
-Among the total 104K QA pairs, approximately half are multiple-choice and the other half are open-ended.
-This entire data procedure consumes about 80,000 GPU hours on H100s.
+Among the total 104K QA pairs, approximately half are multiple-choice and hald are open-ended.
+This entire data procedure consumes about 80,000 H100 GPU hours.
 4 LongVILA Training Pipeline
-As shown in Figure 5, there are two extended training stages in LongVILA-R1, i.e., (1) warm-up for long video reasoning, utilizing 36K data with high-quality CoT for SFT on the MM-SP system [4]; (2) reinforcement learning with dense frames from long videos.
+As shown in Figure 5, there are two extended training stages in LongVILA-R1, i.e., (1) warm-up for long video reasoning, utilizing 36K data with high quality CoT for SFT on the MM-SP system [4]; (2) reinforcement learning with dense frames from long videos.
 4.1 Long Video CoT Supervised Fine-Tuning
-Utilizing 104K high-quality question-reasoning-answer pairs, we apply the data filtering method described in Section 3.1 to select 36K examples for long CoT-SFT, serving as a warm-up phase for subsequent RL.
+Utilizing 104K high quality question-reasoning-answer pairs, we apply the data filtering method described in Section 3.1 to select 36K examples for long CoT-SFT, serving as a warm-up phase for subsequent RL.
 This stage equips the model with fundamental reasoning abilities and instruction-following skills for long video scenarios.
@@
 Then, each GPU locally computes logits for its token slice, prefilling in parallel.
-Table 1: Performance on 6 video benchmarks [43, 44, 14, 13, 45, 20].
-LongVILA-R1-7B outperforms LongVILA-7B across all benchmarks, with varying margins that reflect different levels of reasoning difficulty.
+Table 1: Performance on 6 video benchmarks [43, 44, 14, 13, 45, 20]. LongVILA-R1-7B outperforms LongVILA-7B across all benchmarks, with varying margins.
 Model ActivityNet-QA LongVideoBench PerceptionTest NExT-QA VNBench VideoMME
 test val val mc val w/o sub. w/ sub.
-GPT-4V 57.0 61.3 - - - 59.9 63.3
-GPT-4o - 66.7 - - 64.4 71.9 77.2
+GPT-4o mini - 56.5 - - - 64.8 68.9
+GPT-4o 61.9 66.7 - - 64.4 71.9 77.2
 Gemini-1.5-Pro 57.5 64.0 - - 66.7 75.0 81.3
@@
 LongVILA-7B 59.5 57.1 58.1 80.7 63.0 60.1 65.1
-LongVILA-R1-7B 64.8 58.0 68.9 81.5 75.5 65.0 70.7
+LongVILA-R1-7B 64.8 58.0 68.9 81.5 75.5 65.1 71.1
 Table 2: Performance on LongVideo-Reason-eval set. LongVILA-R1-7B achieves the highest overall score.
@@
   Gemini-1.5-Pro [18]   65.4   81.9   67.8   53.3   69.3
+  LongVILA-7B   58.0   80.2   57.1   46.7   62.7
   LongVILA-R1-7B   68.1   85.7   70.6   53.3   72.0
@@
 Overall Short Medium Long Overall Short Medium Long
+Video-R1-7B [2] 61.4 - - - - - - -
+Apollo-7B [46] 61.1 - - - 63.3 - - -
+LLaVA-Video-7B [12] 63.3 - - - 69.7 - - -
+NVILA-8B-Video [11] 64.2 - - - 70.0 - - -
 Video-LLaVA-7B [47] 39.9 45.3 38.0 36.2 41.6 46.1 40.7 38.1
@@
 VideoChat2-7B [50] 39.5 48.3 37.0 33.2 43.8 52.8 39.4 39.2
-VideoLLaMA2-7B [50] 47.9 56.0 45.4 42.1 50.3 59.4 47.6 43.8
 Chat-Univi-v1.5-7B [51] 40.6 45.7 40.3 35.8 45.9 51.2 44.6 41.8
@@
 LongVA-7B [1] 52.6 61.1 50.4 46.2 54.3 61.1 53.6 47.6
-Video-R1-7B [2] 61.4 - - - - - - -
+VITA-1.5-7B [54] 56.1 67.0 54.2 47.1 58.7 69.9 55.7 50.4
 LongVILA-7B [4] 60.1 69.0 58.3 53.0 65.1 72.9 64.9 57.4
-LongVILA-R1-7B 65.0 76.6 63.6 54.9 70.7 79.0 69.1 64.0
+LongVILA-R1-7B 65.1 76.8 63.2 55.2 71.1 79.2 69.7 64.3
 6 Experimental Results
@@
 LongVILA-R1-7B consistently outperforms LongVILA-7B across all benchmarks, with performance gaps that vary according to the complexity of the reasoning tasks.
-Table 3 presents the general performance of LongVILA-R1, comparing to existing advanced models [46, 47, 48, 49, 50, 51, 52, 53, 1, 2] under comparable model sizes on the Video-MME [20] benchmark.
+Table 3 presents the general performance of LongVILA-R1, comparing to existing advanced models [47, 48, 49, 50, 51, 52, 53, 1, 2, 46, 55, 56, 12, 11, 57] under comparable model sizes on the Video-MME [20] benchmark.
 The LongVILA-R1-7B is tested using 512 video frames as inputs.
-LongVILA-R1-7B achieves leading scores across different video lengths, obtaining scores of 65.0% and 70.7% in the settings without subtitles and with subtitles, respectively.
+LongVILA-R1-7B achieves leading scores across different video lengths, obtaining scores of 65.1% and 71.1% in the settings without subtitles and with subtitles, respectively.
 Table 2 compares the results of our LongVideo-Reason-eval benchmark.
-We also show some qualitative results in the appendix.
 LongVILA-R1-7B model achieves a strong performance with an average score of 72.0%, significantly surpassing Video-R1-7B [2] and slightly outperforms Gemini-1.5-Pro [18].
-Figure 8: Scaling video frames improves long video reasoning performance against non-reasoning baselines.
+Qualitative results are in the appendix.
 6.2 Ablation Study
@@
 Notably, LongVILA-R1-1.5B demonstrates steady performance improvements throughout the scaling process.
-In contrast, LongVILA-1.5B hits a performance bottleneck with 256 frames, and got a degradation on 512 frames, due to its inability to effectively process such dense visual information.
+In contrast, LongVILA-1.5B hits a performance bottleneck with 256 frames, and got a degradation on 512 frames.
 The enhanced reasoning capabilities of LongVILA-R1-1.5B allow it to effectively integrate and infer information from long videos.
+Figure 8: Scaling video frames improves long video reasoning performance against non-reasoning baselines.
 Ablation on pipeline and datasets.
@@
 Leveraging our curated dataset of 52K long video question-reasoning-answer pairs, combined with other open-source video datasets, we adopt a two-stage training process that integrates CoT-SFT and RL.
-LongVILA-R1-7B demonstrates outstanding performance on mainstream video benchmarks, achieving 65.0% and 70.7% on VideoMME without and with subtitles.
-LongVILA-R1 maintains consistent performance improvements, extending the feasible frame count from 16 to 512.
-Notably, our MR-SP leads to a speed up of 2.1$\times$ for long video RL training, supporting hour-level (3,600 frames / 256k tokens) RL training on a single node of 8 A100 GPUs.
-In addition, we release our training system to the public, which supports RL training across multiple modalities (video, text, and audio), various models (including the VILA and Qwen series), and even image and video generation models.
-In the appendix, we discuss RL training curves, border impacts, detailed data generation process, and provide more examples.
-Limitations
+LongVILA-R1-7B demonstrates outstanding performance on mainstream video benchmarks, achieving 65.1% and 71.1% on VideoMME without and with subtitles.
+LongVILA-R1 supports processing up to 8,192 video frames per video, and configurable FPS settings.
+Notably, our MR-SP leads to a speed up of 2.1$\times$ for long video RL training, supporting hour-level (3,600 frames) RL training on a single node of 8 A100 GPUs.
+In addition, we release our training system to the public, which supports RL training across multiple modalities (video, text, and audio), various models (including VILA and Qwen series), and even image and video generation models.
+Appendix A Limitations
 While our RL training system is capable of handling long video frames efficiently (3,600 frames on a single node of A100 GPUs), scaling to significantly longer sequences, more modalities (like include audio for omni VLMs) or large batch sizes would require distributed training across multiple GPU machines.
 This requires more GPU machines, making it less feasible to run on limited hardware resources.
-Appendix A Border Impacts
+Appendix B Border Impacts
 The development of LongVILA-R1 represents a significant advancement in long video reasoning, enabling real world systems to perform sophisticated temporal and compositional understanding across diverse and extended visual contexts.
@@
 By addressing these challenges, the AI community can ensure that advancements in long video reasoning contribute positively to society while mitigating associated risks.
-Appendix B More Demos in Different Reasoning Subsets
+Appendix C More Demos in Different Reasoning Subsets
 In this section, we selected examples of four types of reasoning tasks from the benchmark to evaluate and compare the reasoning processes and answers provided by Gemini-1.5-Pro, Video-R1-7B, and LongVILA-R1-7B.

### v4

@@
 Abstract
-Abstract: We introduce a full-stack framework that scales up reasoning in vision-language models (VLMs) to long videos, leveraging reinforcement learning.
-We address the unique challenges of long video reasoning by integrating three critical components: (1) a large-scale dataset, LongVideo-Reason, comprising 104K long video QA pairs with high-quality reasoning annotations across diverse domains such as sports, games, and vlogs; (2) a two-stage training pipeline that extends VLMs with chain-of-thought supervised fine-tuning (CoT-SFT) and reinforcement learning (RL); and (3) a training infrastructure for long video RL, named Multi-modal Reinforcement Sequence Parallelism (MR-SP), which incorporates sequence parallelism and a vLLM-based engine tailored for long video, using cached video embeddings for efficient rollout and prefilling.
-In our experiments, LongVILA-R1-7B achieves strong performance on video benchmarks, reaching 65.1% and 71.1% accuracy on VideoMME without and with subtitles, respectively, and consistently outperforming LongVILA-7B across multiple benchmarks.
-Moreover, LongVILA-R1-7B supports processing up to 8,192 video frames per video, and configurable FPS settings.
-Notably, our MR-SP system achieves up to 2.1$\times$ speedup on long video RL training.
-In addition, we release our training system for public availability that supports RL training on various modalities (video, text, and audio), various models (VILA and Qwen series), and even image and video generation models.
-On a single A100 node (8 GPUs), it supports RL training on hour-long videos (e.g., 3,600 frames).
-Code and models are available at https://github.com/NVlabs/Long-RL and https://huggingface.co/Efficient-Large-Model/LongVILA-R1-7B.
+[…]
+00footnotetext: Equal contribution    Core contribution
 1 Introduction
-Understanding long videos requires more than simple recognition—it demands reasoning from temporal, spatial, goal-oriented, and narrative perspectives [1].
+Understanding long videos requires more than simple recognition—it demands reasoning from temporal, spatial, goal-oriented, and narrative perspectives zhang2024long .
 As illustrated in Figure 1, answering high-level questions often hinges on a model’s ability to integrate clues distributed across time, infer hidden goals or strategies, track entities spatially, and comprehend the evolving plot.
@@
 These examples underscore that reasoning is indispensable for long video understanding that goes beyond recognition alone.
-Figure 1: Training efficiency comparison with MR-SP (SP degree=4) on Qwen2.5-VL-7B and LongVILA-R1-7B and a single node 8$\times$ A100 GPUs.
-It achieves 2.1$\times$ speed-up and avoids GPU OOM issue on long frames.
+Despite the clear importance of reasoning in long video understanding, enabling such capabilities in long video VLMs poses significant challenges feng2025video ; weng2024longvlm ; chen2024longvila ; guo2025fila ; shen2025long .
+First, the collection of high quality long video reasoning datasets is inherently difficult.
+Unlike domains such as math or code reasoning, where structured supervision and benchmarks are readily available shao2024deepseekmath ; liu2024codemind , long video reasoning requires annotating complex temporal dynamics, goals, spatial relations, and narrative elements—often across minutes or hours of footage han2024videoespresso .
+This process is labor-intensive and subjective, making large-scale dataset construction slow and costly.
+Second, the RL training framework for long videos is challenging.
+Reinforcement learning, a common strategy for aligning models with complex reasoning objectives, is computationally expensive and sample-inefficient sheng2024hybridflow .
+When applied to long videos, RL becomes even more burdensome due to the extended video frames, requiring more memory and longer rollout runtime.
+These challenges jointly hinder the development of effective long video VLMs with strong reasoning capabilities.
 Figure 1: Examples of LongVILA-R1.
@@
 Notably, the spatial tracking video involves a relatively complex dynamic moving, for which the model fails to achieve accurate reasoning until the number of input video frames increases to 128.
-Despite the clear importance of reasoning in long video understanding, enabling such capabilities in long video VLMs poses significant challenges [2, 3, 4, 5, 6].
-First, the collection of high quality long video reasoning datasets is inherently difficult.
-Unlike domains such as math or code reasoning, where structured supervision and benchmarks are readily available [7, 8], long video reasoning requires annotating complex temporal dynamics, goals, spatial relations, and narrative elements—often across minutes or hours of footage [9].
-This process is labor-intensive and subjective, making large-scale dataset construction slow and costly.
-Second, the RL training framework for long videos is challenging.
-Reinforcement learning, a common strategy for aligning models with complex reasoning objectives, is computationally expensive and sample-inefficient [10].
-When applied to long videos, RL becomes even more burdensome due to the extended video frames, requiring more memory and longer rollout runtime.
-These challenges jointly hinder the development of effective long video VLMs with strong reasoning capabilities.
 In this work, we introduce LongVILA-R1, a comprehensive framework exploring the reasoning capabilities for long video understandings.
 Firstly, we strategically construct a high quality dataset with CoT annotations for long video reasoning, named LongVideo-Reason.
-Leveraging a powerful VLM (NVILA-8B) [11] and a leading open-source reasoning LLM, we develop a dataset comprising 104K high quality Question-Reasoning-Answer pairs for long videos.
-We use 36K high quality samples for Long-CoT-SFT to initialize the model’s reasoning and instruction-following abilities, and 68K samples with an additional 102K video data [12, 13, 14, 15, 16] for reinforcement learning.
+Leveraging a powerful VLM (NVILA-8B) liu2024nvila and a leading open-source reasoning LLM, we develop a dataset comprising 104K high quality Question-Reasoning-Answer pairs for long videos.
+We use 36K high quality samples for Long-CoT-SFT to initialize the model’s reasoning and instruction-following abilities, and 68K samples with an additional 102K video data llava-video ; next-qa ; perpceptiontest ; clevr ; star for reinforcement learning.
 This two-stage training combines high quality reasoning annotations with reinforcement learning, enabling LongVILA-R1 to achieve superior and generalized video reasoning.
 We also manually curate a balanced set of 1K long video samples to build a new benchmark, LongVideo-Reason-eval, that evaluates performance from four perspectives: Temporal, Goal and Purpose, Spatial, and Plot and Narrative, for a comprehensive assessment.
-Figure 3: Data Distribution of LongVideo-Reason and total data in the LongVILA-R1 training framework.
-LongVideo-Reason comprises a total of 18K videos and 104K QAs with reasoning annotations.
-Additionally, we also include additional 102K QAs from existing works [12, 13, 14, 15, 16] for RL training.
-Secondly, we propose a training framework for VLMs to advance long video reasoning.
-As illustrated in Figure 5, this framework incorporates two stages, i.e., Stage-1: Long CoT-SFT, and Stage-2: RL for long video reasoning.
-To address the unique challenges of long video RL, including massive visual embeddings, heavy rollouts, and long-context LLM prefilling, we develop an efficient and scalable solution, referred to as Multi-modal Reinforcement Sequence Parallelism (MR-SP).
-It incorporates a vLLM engine [17] tailored for LongVILA and a caching scheme for video embeddings.
-The MR-SP system alleviates the problem of intensive memory and facilitates RL training of long video VLMs.
-As shown in Figure 1, our MR-SP system achieves up to 2.1$\times$ speedup on 512-frame video RL training on 7B models and enables longer training frames without out-of-memory (OOM).
+[…]
+[…]
 In our experiments, LongVILA-R1-7B demonstrates strong performance on video benchmarks, achieving 65.1% and 71.1% accuracy on VideoMME without and with subtitles, respectively.
@@
 Additionally, LongVILA-R1 supports processing up to 8,192 video frames per video, and configurable FPS settings.
-On our LongVideo-Reason-eval benchmark, LongVILA-R1-7B achieves an average accuracy of 72.0%, surpassing open-source models like Video-R1-7B[2], as well as proprietary models such as Gemini-1.5-Pro[18].
+On our LongVideo-Reason-eval benchmark, LongVILA‑R1‑7B achieves an average accuracy of 72.0%, surpassing Video‑R1‑7Bfeng2025video , as well as proprietary models such as Gemini‑1.5‑Proteam2024gemini .
+2 Related Work
+Multi-modal reasoning models.
+The field of multi-modal reasoning has advanced significantly, particularly in Vision-Language Models (VLMs).
+GPT-4o gpt4o improves visual understanding through enhanced reasoning, while Gemini-1.5-Pro team2024gemini extends context length to 1 million tokens, achieving state-of-the-art performance in VideoMME video-mme .
+Following the substantial progress of architecture and training algorithms in open-source VLMs zhu2025internvl3 ; liu2024nvila ; yang2024qwen2 ; lin2024vila ; vila-hd , multi-modal reasoning has been further explored in works including LMM-R1 peng2025lmm which employs a two-stage training strategy, Vision-R1 huang2025vision that addresses post-cold-start overthinking, and Video-R1 feng2025video which enhances RL for video via T-GRPO in 16 frames.
+However, these approaches primarily focus on single images or short videos, and long video reasoning still poses great challenges.
+Figure 3: Data Distribution of LongVideo-Reason and total data in the LongVILA-R1 training framework.
+LongVideo-Reason comprises a total of 18K videos and 104K QAs with reasoning annotations.
+Additionally, we include 102K QAs from existing works llava-video ; next-qa ; perpceptiontest ; clevr ; star .
+Sequence parallelism.
+Training with long contexts often exceeds the memory capacity of a single device longlora , necessitating efficient distribution strategies.
+Sequence parallelism (SP) has become a widely adopted solution  liu2023ring ; jacobs2023deepspeed ; InternEvo ; USP ; LoongTrain .
+For example, ring-based systems like LightSeq li2023lightseq and Ring Attention liu2023ring use point-to-point (P2P) communication, while DeepSpeed-Ulysses jacobs2023deepspeed employs all-to-all (A2A) primitives to optimize attention computations.
+Additionally, USP and LoongTrain  USP ; LoongTrain were introduced to integrate Ring-style SP and Ulysses SP.
+LongVILA chen2024longvila further proposed multi-modal SP (MM-SP), enabling vision-language models to handle long-context inputs.
+However, multi-modal reinforcement learning introduces additional challenges, as it requires extensive sampling from long, mixed-token sequences sheng2024hybridflow , particularly in complex group optimization tasks shao2024deepseekmath .
+RL frameworks for LLMs/VLMs.
+[…]
+3 LongVideo-Reason Data Construction
+3.1 Overview of Data Curation
+We first curate 18K long videos from the Shot2Story dataset han2023shot2story20k (Figure 3, left).
+We further incorporated 2k additional 4K-resolution videos spanning scenarios such as autonomous driving, video games, household robotics, and wildlife.
+We then apply a high quality automated annotation pipeline for CoT as detailed in Section 3.2, and end up with a total of 104K Question-Reasoning-Answer pairs where each sample, based on the type of question it is reasoning about, can be categorized into Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning (Figure 3, middle).
+This dataset is designed to support various types of long-video reasoning tasks comprehensively.
+Given the sensitivity of GRPO to batch sampling team2025kimi ; peng2025lmm , a data filtering approach is adopted.
+Specifically, we employ a test-scaling method where LongVILA chen2024longvila performs inference 10 times on the original datasets.
+Questions consistently answered correctly or incorrectly are labeled as easy or hard while those inducing diverse predictions are labeled as medium.
+We filter out too easy and too hard questions, and use remaining questions for training.
+The reason is that GRPO expects different rollouts of each sample to be diverse in order to have meaningful advantages, and the gradient vanishes if all the rollouts predict correct or incorrect answers.
+The COT-SFT subset (36K) features high quality CoT reasoning processes formatted in a standard <think></think><answer></answer> structure, providing abundant resources for warm-up training during Stage 1 of the model’s reasoning capabilities.
+Meanwhile, the RL subset contains 68K challenging long-video Q&A, which is leveraged in Stage 2 for scaling reasoning through reinforcement learning.
+To further enhance RL scaling, we incorporate an additional 102K high quality open-source videos (Figure 3, right) from other datasets llava-video ; next-qa ; perpceptiontest ; clevr ; star .
+This combination improves the model’s generalization.
 Figure 4: Data generation process for the LongVideo-Reason dataset.
@@
 Finally, the reasoning annotations are reformatted for conciseness and alignment with video details.
-We present a more detailed figure of data generation process in Figure 9 in the appendix.
-2 Related Work
-Multi-modal reasoning models.
-The field of multi-modal reasoning has advanced significantly, particularly in Vision-Language Models (VLMs).
-GPT-4o [19] improves visual understanding through enhanced reasoning, while Gemini-1.5-Pro [18] extends context length to 1 million tokens, achieving state-of-the-art performance in VideoMME [20].
-Following the substantial progress of architecture and training algorithms in open-source VLMs [21, 11, 22, 23, 24], multi-modal reasoning has been further explored in works including LMM-R1 [25] which employs a two-stage training strategy, Vision-R1 [26] that addresses post-cold-start overthinking, and Video-R1 [2] which enhances RL for video via T-GRPO in 16 frames.
-However, these approaches primarily focus on single images or short videos, and long video reasoning still poses great challenges.
-Sequence parallelism.
-Training with long contexts often exceeds the memory capacity of a single device [27], necessitating efficient distribution strategies.
-Sequence parallelism (SP) has become a widely adopted solution  [28, 29, 30, 31, 32].
-For example, ring-based systems like LightSeq [33] and Ring Attention [28] use point-to-point (P2P) communication, while DeepSpeed-Ulysses [29] employs all-to-all (A2A) primitives to optimize attention computations.
-Additionally, USP and LoongTrain  [31, 32] were introduced to integrate Ring-style SP and Ulysses SP.
-LongVILA [4] further proposed multi-modal SP (MM-SP), enabling vision-language models to handle long-context inputs.
-However, multi-modal reinforcement learning introduces additional challenges, as it requires extensive sampling from long, mixed-token sequences [10], particularly in complex group optimization tasks [7].
-RL frameworks for LLMs/VLMs.
-Reinforcement Learning (RL) has become a key strategy for enhancing Large Language Models (LLMs), particularly through Reinforcement Learning with Human Feedback (RLHF) [34] or Direct Preference Optimization [35], which aligns model outputs with human preferences.
-Recent advancements demonstrate that RL significantly improves LLM reasoning abilities.
-For example, DeepSeek-R1 [36] utilizes the Group Relative Policy Optimization (GRPO) algorithm [37], integrating group-based sampling and rule-based rewards.
-On the other hand, RL [38] poses a unique challenge of heavy computational cost, especially in multi-modal settings.
-To address this problem, HybridFlow [10] is introduced, leveraging Ray [39] for efficient data flow and vLLM [17] for faster sampling.
-Nevertheless, this remains a bottleneck when processing long video sequences, with group-based sampling constrained by the high computational cost of long-context sampling and visual encoding.
-In this work, we propose MR-SP that provides up to 2.1$\times$ speedup for long video RL training.
-3 LongVideo-Reason Data Construction
-3.1 Overview of Data Curation
-We first curate 18K long videos from the Shot2Story dataset [40] (Figure 3, left).
-We further incorporated 2k additional 4K-resolution videos spanning scenarios such as autonomous driving, video games, household robotics, and wildlife.
-We then apply a high quality automated annotation pipeline for CoT as detailed in Section 3.2, and end up with a total of 104K Question-Reasoning-Answer pairs where each sample, based on the type of question it is reasoning about, can be categorized into Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning (Figure 3, middle).
-This dataset is designed to support various types of long-video reasoning tasks comprehensively.
-Given the sensitivity of GRPO to batch sampling [41, 25], a data filtering approach is adopted.
-Specifically, we employ a test-scaling method where LongVILA [4] performs inference 10 times on the original datasets.
-Questions consistently answered correctly or incorrectly are labeled as easy or hard while those inducing diverse predictions are labeled as medium.
-We filter out too easy and too hard questions, and use remaining questions for training.
-The reason is that GRPO expects different rollouts of each sample to be diverse in order to have meaningful advantages, and the gradient vanishes if all the rollouts predict correct or incorrect answers.
-The COT-SFT subset (36K) features high quality CoT reasoning processes formatted in a standard <think></think><answer></answer> structure, providing abundant resources for warm-up training during Stage 1 of the model’s reasoning capabilities.
-Meanwhile, the RL subset contains 68K challenging long-video Q&A, which is leveraged in Stage 2 for scaling reasoning through reinforcement learning.
-To further enhance RL scaling, we incorporate an additional 102K high quality open-source videos (Figure 3, right) from other datasets [12, 13, 14, 15, 16].
-This combination improves the model’s generalization.
-Figure 5: The LongVILA-R1 training pipeline.
-LongVILA-R1 builds upon the base training pipeline for LongVILA.
-MM-SP is further employed for SFT on long video understanding tasks with long CoT.
-Then, reinforcement scaling learning is conducted through Multi-modal Reinforcement Sequential Parallelism (MR-SP).
-Figure 6: LongVILA-R1 RL training framework.
-The framework integrates multi-modal reinforcement sequence parallelism (MR-SP) for scalable video frame encoding and LLM prefilling.
-RL utilizes a vLLM-based engine with cached video embeddings, tailored for LongVILA rollout.
-Rewards for accuracy and format guide policy optimization.
+We present a more detailed figure of data generation process in Figure 9.
 3.2 Long-Video Reasoning Generation
 We introduce an automated annotation pipeline (Figure 4) that generates high quality Question-Reasoning-Answer pairs from long videos.
-This process begins by segmenting videos into short clips (~10 seconds each), each of which is annotated using the NVILA-8B [11] model to provide descriptive captions.
-For the spatial reasoning category, We employed VILA-HD [24] to generate object bounding boxes in video frames and constructed spatial reasoning QAs based on these boxes and corresponding captions.
-Leveraging the breakthrough in text-based reasoning, we then deploy a leading open-source reasoning LLM, DeepSeek-R1-671B [36], provide the captions of all the clips in each video, and prompt it to generate diverse types of Question-Reasoning-Answer pairs that involve reasoning over the content across the whole video.
+This process begins by segmenting videos into short clips (~10 seconds each), each of which is annotated using the NVILA-8B liu2024nvila model to provide descriptive captions.
+For the spatial reasoning category, We employed VILA-HD vila-hd to generate object bounding boxes in video frames and constructed spatial reasoning QAs based on these boxes and corresponding captions.
+Leveraging the breakthrough in text-based reasoning, we then deploy a leading open-source reasoning LLM, DeepSeek-R1-671B guo2025deepseek , provide the captions of all the clips in each video, and prompt it to generate diverse types of Question-Reasoning-Answer pairs that involve reasoning over the content across the whole video.
 Specifically, we design four types of prompts to encourage the LLM to generate a Question-Reasoning-Answer pair that focuses on one of the four types of reasoning: Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning.
@@
 4 LongVILA Training Pipeline
-As shown in Figure 5, there are two extended training stages in LongVILA-R1, i.e., (1) warm-up for long video reasoning, utilizing 36K data with high quality CoT for SFT on the MM-SP system [4]; (2) reinforcement learning with dense frames from long videos.
+As shown in Figure 5, there are two extended training stages in LongVILA-R1, i.e., (1) warm-up for long video reasoning, utilizing 36K data with high quality CoT for SFT on the MM-SP system chen2024longvila ; (2) reinforcement learning with dense frames from long videos.
 4.1 Long Video CoT Supervised Fine-Tuning
@@
 This stage equips the model with fundamental reasoning abilities and instruction-following skills for long video scenarios.
-To efficiently perform SFT on hundreds of frames, we adopt the MM-SP [4] training system from LongVILA.
-As demonstrated in Section 6.2, SFT solely on our LongVideo-Reason dataset also effectively improves the model with foundational reasoning capabilities.
-Figure 7: The workflow of multi-modal reinforcement sequential parallel (MR-SP).
-To accommodate multi-modal input in reinforcement learning, we develop a custom sharding strategy to ensure balanced workload distribution and compatibility with SP communication.
-Efficient video embedding reuse and vLLM rollout acceleration strategies are implemented, while meeting the demands of policy model prefilling for dense video frames.
+To efficiently perform SFT on hundreds of frames, we adopt the MM-SP chen2024longvila training system from LongVILA.
+As demonstrated in Section 6.2, SFT solely on our LongVideo-Reason dataset also effectively improves the model with reasoning capabilities.
 4.2 GRPO for Long Video
-Building on the advancements of the GRPO [7] algorithm and prior explorations of multi-modal reasoning training [2, 25], we adhere to the standard GRPO framework to train our model.
+Building on the advancements of the GRPO shao2024deepseekmath algorithm and prior explorations of multi-modal reasoning training feng2025video ; peng2025lmm , we adhere to the standard GRPO framework to train our model.
 For each given question $q$, the policy model generates a group of candidate responses $\{o_{1},o_{2},...,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$, accompanied by their corresponding rewards $\{r_{1},r_{2},...,r_{G}\}$, which are computed based on rule-based reward functions (format/accuracy).
@@
 $A_{i}=\frac{r_{i}-\mathrm{mean}(\{r_{1},r_{2},...,r_{G}\})}{\mathrm{std}(\{r_{1},r_{2},...,r_{G}\})}$ (2)
+Figure 5: The LongVILA-R1 training pipeline.
+LongVILA-R1 builds upon the base training pipeline for LongVILA.
+$M$-SP is further employed for SFT on long video understanding tasks with long CoT.
+Then, reinforcement scaling learning is conducted through Multi-modal Reinforcement Sequential Parallelism (MR-SP).
+Figure 6: LongVILA-R1 RL training framework.
+The framework integrates multi-modal reinforcement sequence parallelism (MR-SP) for scalable video frame encoding and LLM prefilling.
+RL utilizes a vLLM-based engine with cached video embeddings, tailored for LongVILA rollout.
+Rewards for accuracy and format guide policy optimization.
 However, RL for long videos presents significant challenges due to the high computational demands of processing hundreds to thousands of frames.
@@
 To address this, we develop the MR-SP framework (Section 5), which efficiently scales reinforcement learning for long-context video reasoning.
-Considering the sensitivity of GRPO to sampling during training [41], we use the 68K filtered data for reinforcement learning as described in Section 3.1.
-Additionally, an extra 102K samples from other datasets [12, 13, 14, 15, 16] are incorporated to scale up the RL.
+Considering the sensitivity of GRPO to sampling during training team2025kimi , we use the 68K filtered data for reinforcement learning as described in Section 3.1.
+Additionally, an extra 102K samples from other datasets llava-video ; next-qa ; perpceptiontest ; clevr ; star are incorporated to scale up the RL.
 This approach aims to guide the model in freely exploring and developing more effective and generalized reasoning strategies.
 5 Multi-modal Reinforcement SP
-Existing RL frameworks for VLMs, such as R1-V [38] and EasyR1 [42], are not designed for long videos which present unique challenges due to their high token volume.
+Existing RL frameworks for VLMs, such as R1-V chen2025r1v and EasyR1 EasyR1 , are not designed for long videos which present unique challenges due to their high token volume.
 To address this, we introduce Multi-modal Reinforcement Sequence Parallelism (MR-SP), a framework for efficient RL training on long videos.
 MR-SP leverages sequence parallelism in both rollout and pre-filling stages, enabling long videos in RL, with reduced overhead.
+We show the training curve with MR-SP in Figure 8.
 5.1 Stage 1 - Rollout with Paralleled Encoding
@@
 This strategy distributes the encoding workload, allowing the system to handle significantly longer videos by leveraging more GPUs, while avoiding the risk of GPU memory overflow.
-The parallel encoding scheme ensures balanced utilization of the vision towers and enables scalable long-video processing that would otherwise be infeasible on a single device.
+The parallel encoding scheme ensures balanced utilization of the vision towers and enables scalable long-video processing that would otherwise be infeasible.
 After the video embeddings are globally gathered, they are reused for downstream usage throughout the RL pipeline.
@@
 By caching and reusing the gathered embeddings, MR-SP eliminates this redundancy and significantly accelerates training.
+Figure 7: The workflow of multi-modal reinforcement sequential parallel (MR-SP).
+To accommodate multi-modal input in reinforcement learning, we develop a custom sharding strategy to ensure balanced workload distribution and compatibility with SP communication.
+Efficient video embedding reuse and vLLM rollout acceleration strategies are implemented, while meeting the demands of policy model prefilling for dense video frames.
+Figure 8: Reward curve of our LongVILA-R1-7B with MR-SP. During training, LongVILA-R1-7B exhibits stable reward improvements across overall, format, and accuracy dimensions.
 5.2 Stage 2 - Prefilling with Sequence Parallelism
@@
 Then, each GPU locally computes logits for its token slice, prefilling in parallel.
-Table 1: Performance on 6 video benchmarks [43, 44, 14, 13, 45, 20]. LongVILA-R1-7B outperforms LongVILA-7B across all benchmarks, with varying margins.
-Model ActivityNet-QA LongVideoBench PerceptionTest NExT-QA VNBench VideoMME
+6 Experimental Results
+6.1 Main Results
+Table 1 shows the performance comparison on 6 video benchmarks (activity-qa, ; longvideobench, ; perpceptiontest, ; next-qa, ; vnbench, ; video-mme, ).
+LongVILA-R1-7B consistently outperforms LongVILA-7B across all benchmarks, with performance gaps that vary according to the complexity of the reasoning tasks.
+Table 3 presents the general performance of LongVILA-R1, comparing to existing advanced models lin2023video ; zhang2024beyond ; chen2024sharegpt4video ; li2023videochat ; jin2024chat ; liu2024kangaroo ; wang2024retake ; zhang2024long ; feng2025video ; apollo ; qwen2-vl ; internvl2.5 ; llava-video ; liu2024nvila ; videollama3 under comparable model sizes on the Video-MME video-mme benchmark.
+The LongVILA-R1-7B is tested using 512 video frames as inputs.
+LongVILA-R1-7B achieves leading scores across different video lengths, obtaining scores of 65.1% and 71.1% in the settings without subtitles and with subtitles, respectively.
+Table 2 compares the results of our LongVideo-Reason-eval benchmark.
+LongVILA-R1-7B model achieves a strong performance with an average score of 72.0%, surpassing Video-R1-7B feng2025video and slightly outperforms Gemini-1.5-Pro team2024gemini .
+Qualitative results are in the appendix.
+6.2 Ablation Study
+Scaling video frames.
+The reasoning capability of LongVILA-R1 scales consistently with the number of input video frames.
+Specifically, Table 5 illustrates the performance of LongVILA-1.5B (grey line) and LongVILA-1.5B-R1 (red line) on the long-video reasoning benchmark under varying frame inputs.
+With only 16 input frames, LongVILA-R1-1.5B is close to LongVILA-1.5B.
+However, as the number of frames increases to 512, LongVILA-R1-1.5B consistently outperforms and eventually achieves a score of 64.3%.
+Notably, LongVILA-R1-1.5B demonstrates steady performance improvements throughout the scaling process.
+In contrast, LongVILA-1.5B hits a performance bottleneck with 256 frames, and got a degradation on 512 frames.
+The enhanced reasoning capabilities of LongVILA-R1-1.5B allow it to effectively integrate and infer information from long videos.
+Table 1: Performance on ActivityNet-QA (activity-qa, ), LongVideoBench (longvideobench, ), PerceptionTest (perpceptiontest, ), NExT-QA (next-qa, ), VNBench (vnbench, ), and VideoMME (video-mme, ).
+LongVILA-R1-7B outperforms LongVILA-7B across all these benchmarks, with varying margins.
+Model Act.Net-QA L.V.Bench Per.Test NExT-QA VNBench VideoMME
 test val val mc val w/o sub. w/ sub.
@@
 LongVILA-R1-7B 64.8 58.0 68.9 81.5 75.5 65.1 71.1
-Table 2: Performance on LongVideo-Reason-eval set. LongVILA-R1-7B achieves the highest overall score.
+Table 2: Performance on LongVideo-Reason-eval. LongVILA-R1-7B achieves a strong overall score.
   Model   Temporal   Goal   Plot   Spatial   Overall
-  Video-R1-7B [2]   61.4   85.0   62.0   58.5   68.1
-  Gemini-1.5-Pro [18]   65.4   81.9   67.8   53.3   69.3
+  Video-R1-7B feng2025video   61.4   85.0   62.0   58.5   68.1
+  Gemini-1.5-Pro team2024gemini   65.4   81.9   67.8   53.3   69.3
   LongVILA-7B   58.0   80.2   57.1   46.7   62.7
   LongVILA-R1-7B   68.1   85.7   70.6   53.3   72.0
-Table 3: Performance comparison on VideoMME [20] benchmark in details.
+Table 3: Performance comparison on VideoMME (video-mme, ) benchmark in details.
 Model w/o subtitle w subtitle
 Overall Short Medium Long Overall Short Medium Long
-Video-R1-7B [2] 61.4 - - - - - - -
-Apollo-7B [46] 61.1 - - - 63.3 - - -
-LLaVA-Video-7B [12] 63.3 - - - 69.7 - - -
-NVILA-8B-Video [11] 64.2 - - - 70.0 - - -
-Video-LLaVA-7B [47] 39.9 45.3 38.0 36.2 41.6 46.1 40.7 38.1
-SliME-8B [48] 45.3 53.3 55.4 39.8 47.2 55.4 44.4 41.7
-ShareGPT4Video-8B [49] 39.9 48.3 36.3 35.0 43.6 53.6 39.3 37.9
-VideoChat2-7B [50] 39.5 48.3 37.0 33.2 43.8 52.8 39.4 39.2
-Chat-Univi-v1.5-7B [51] 40.6 45.7 40.3 35.8 45.9 51.2 44.6 41.8
-Kangaroo-8B [52] 56.0 66.1 55.3 46.7 57.6 68.0 55.4 49.3
-ShareGemini-7B [53] 43.2 49.1 41.3 39.1 47.9 49.1 47.3 43.4
-LongVA-7B [1] 52.6 61.1 50.4 46.2 54.3 61.1 53.6 47.6
-VITA-1.5-7B [54] 56.1 67.0 54.2 47.1 58.7 69.9 55.7 50.4
-LongVILA-7B [4] 60.1 69.0 58.3 53.0 65.1 72.9 64.9 57.4
+Video-R1-7B feng2025video 61.4 - - - - - - -
+Apollo-7B apollo 61.1 - - - 63.3 - - -
+LLaVA-Video-7B llava-video 63.3 - - - 69.7 - - -
+NVILA-8B-Video liu2024nvila 64.2 - - - 70.0 - - -
+Video-LLaVA-7B lin2023video 39.9 45.3 38.0 36.2 41.6 46.1 40.7 38.1
+SliME-8B zhang2024beyond 45.3 53.3 55.4 39.8 47.2 55.4 44.4 41.7
+ShareGPT4Video-8B chen2024sharegpt4video 39.9 48.3 36.3 35.0 43.6 53.6 39.3 37.9
+VideoChat2-7B li2023videochat 39.5 48.3 37.0 33.2 43.8 52.8 39.4 39.2
+Chat-Univi-v1.5-7B jin2024chat 40.6 45.7 40.3 35.8 45.9 51.2 44.6 41.8
+Kangaroo-8B liu2024kangaroo 56.0 66.1 55.3 46.7 57.6 68.0 55.4 49.3
+ShareGemini-7B wang2024retake 43.2 49.1 41.3 39.1 47.9 49.1 47.3 43.4
+LongVA-7B zhang2024long 52.6 61.1 50.4 46.2 54.3 61.1 53.6 47.6
+VITA-1.5-7B vita-1.5 56.1 67.0 54.2 47.1 58.7 69.9 55.7 50.4
+LongVILA-7B chen2024longvila 60.1 69.0 58.3 53.0 65.1 72.9 64.9 57.4
 LongVILA-R1-7B 65.1 76.8 63.2 55.2 71.1 79.2 69.7 64.3
-6 Experimental Results
-6.1 Main Results
-Table 1 shows the performance comparison on 6 video benchmarks [43, 44, 14, 13, 45, 20].
-LongVILA-R1-7B consistently outperforms LongVILA-7B across all benchmarks, with performance gaps that vary according to the complexity of the reasoning tasks.
-Table 3 presents the general performance of LongVILA-R1, comparing to existing advanced models [47, 48, 49, 50, 51, 52, 53, 1, 2, 46, 55, 56, 12, 11, 57] under comparable model sizes on the Video-MME [20] benchmark.
-The LongVILA-R1-7B is tested using 512 video frames as inputs.
-LongVILA-R1-7B achieves leading scores across different video lengths, obtaining scores of 65.1% and 71.1% in the settings without subtitles and with subtitles, respectively.
-Table 2 compares the results of our LongVideo-Reason-eval benchmark.
-LongVILA-R1-7B model achieves a strong performance with an average score of 72.0%, significantly surpassing Video-R1-7B [2] and slightly outperforms Gemini-1.5-Pro [18].
-Qualitative results are in the appendix.
-6.2 Ablation Study
-Scaling video frames.
-The reasoning capability of LongVILA-R1 scales consistently with the number of input video frames.
-Specifically, Figure 8 illustrates the performance of LongVILA-1.5B (grey line) and LongVILA-1.5B-R1 (red line) on the long-video reasoning benchmark under varying frame inputs.
-With only 16 input frames, LongVILA-R1-1.5B is close to LongVILA-1.5B.
-However, as the number of frames increases to 512, LongVILA-R1-1.5B consistently outperforms and eventually achieves a score of 64.3%.
-Notably, LongVILA-R1-1.5B demonstrates steady performance improvements throughout the scaling process.
-In contrast, LongVILA-1.5B hits a performance bottleneck with 256 frames, and got a degradation on 512 frames.
-The enhanced reasoning capabilities of LongVILA-R1-1.5B allow it to effectively integrate and infer information from long videos.
-Figure 8: Scaling video frames improves long video reasoning performance against non-reasoning baselines.
 Ablation on pipeline and datasets.
@@
 The accuracies are evaluated on LongVideo-Reason-eval.
-✗  means skipping this stage, ✓  means training this stage with our datasets, and O means training this stage with other datasets [12, 13, 14, 15, 16].
+✗  means skipping this stage, ✓  means training this stage with our datasets, and O means training this stage with other datasets llava-video ; next-qa ; perpceptiontest ; clevr ; star .
 Our CoT-SFT dataset results in a better performance than that of other datasets.
@@
 We measure the forward time for each training step.
-The results are obtained after 10 warming up iterations and averaged over 5 iterations to minimize variance.
+The results are obtained after 10 warming up iterations and averaged over 5 iterations.
 We use LongVILA-7B-R1 model with training batch size as 1 per GPU and rollout number as 5.
@@
 In contrast, the full MR-SP system consistently reduces runtime, achieving up to a 2.1× speedup at 512 frames and scaling efficiently to 1024 frames without OOM, highlighting the benefit of combining sequence reuse and sequence parallelism for long video RL training.
-Table 4: Ablations on training pipeline and datasets.
+Table 4: Ablations on frames and reasoning (R.).
+Frames 16 32 64 128 256 512
+w/o R. 55.7 56.4 58.1 60.5 60.7 60.2
+w/ R. 55.9 56.7 61.9 62.6 64.1 64.3
+Table 5: Ablations on CoT-SFT and RL.
 CoT-SFT O O
@@
 7 Conclusion
-We present a comprehensive framework designed to fully scale VLMs for reasoning over long videos.
-LongVILA-R1 encompasses a meticulously constructed large-scale dataset, LongVideo-Reason, and a parallelized training framework, MR-SP.
-Leveraging our curated dataset of 52K long video question-reasoning-answer pairs, combined with other open-source video datasets, we adopt a two-stage training process that integrates CoT-SFT and RL.
-LongVILA-R1-7B demonstrates outstanding performance on mainstream video benchmarks, achieving 65.1% and 71.1% on VideoMME without and with subtitles.
-LongVILA-R1 supports processing up to 8,192 video frames per video, and configurable FPS settings.
-Notably, our MR-SP leads to a speed up of 2.1$\times$ for long video RL training, supporting hour-level (3,600 frames) RL training on a single node of 8 A100 GPUs.
-In addition, we release our training system to the public, which supports RL training across multiple modalities (video, text, and audio), various models (including VILA and Qwen series), and even image and video generation models.
-Appendix A Limitations
-While our RL training system is capable of handling long video frames efficiently (3,600 frames on a single node of A100 GPUs), scaling to significantly longer sequences, more modalities (like include audio for omni VLMs) or large batch sizes would require distributed training across multiple GPU machines.
-This requires more GPU machines, making it less feasible to run on limited hardware resources.
-Appendix B Border Impacts
+[…]
+Figure 9: Detailed data generation process for the LongVideo-Reason dataset, supplementing Figure 4.
+Limitations
+While our RL training system is capable of handling long video frames efficiently (3,600 frames on a single node of A100 GPUs), scaling to significantly longer sequences, more modalities (like include audio for omni VLMs) or large batch sizes would require distributed training across multiple GPUs.
+This requires more GPUs, making it less feasible to run on limited resources.
+8 Border Impacts
 The development of LongVILA-R1 represents a significant advancement in long video reasoning, enabling real world systems to perform sophisticated temporal and compositional understanding across diverse and extended visual contexts.
 This progress holds transformative potential across multiple domains, including embodied AI, robotics, autonomous systems, and AR/VR applications.
-By equipping vision-language models (VLMs) with the ability to process and reason over long-duration video data, LongVILA-R1 lays the foundation for AI systems capable of understanding event sequences, tracking object persistence and transformations, and inferring causal and physical relationships over extended frames.
+By equipping vision-language models (VLMs) with the ability to process and reason over long-duration video data, LongVILA-R1 lays the foundation for AI systems capable of understanding event sequences, and inferring causal and physical relationships over extended frames.
 Long video reasoning technologies powered by LongVILA-R1 can significantly enhance embodied AI and robotics by enabling agents to sustain coherent, long-term understanding of their environment.
@@
 Furthermore, such systems could enhance sports analytics, and other areas requiring nuanced temporal reasoning.
+Figure 9 shows the the data-generation where our supervision is built from grouped short clips that form long videos, per-clip captions, and LLM-generated Q&A with Long-CoT that is formatted as multiple-choice reasoning rather than identity labels or free-form attribution.
+This design choice intentionally steers the supervision signal toward temporal and causal evidence (what happens, when, and why) and away from sensitive attributes (who a person is), thereby reducing privacy and profiling risks at the data layer.
+Concretely, we derive questions from captions and temporally grounded observations instead of collecting personal metadata; we avoid tasks that require demographic inference, face recognition, or persistent identity tracking; and we release structured Q&A rather than user-originating personal content.
+In this way, the dataset construction supports the intended societal benefits of long-video reasoning while mitigating risks such as privacy leakage, surveillance-like use, and stereotype amplification.
 In conclusion, LongVILA-R1 demonstrates the potential of long video reasoning to drive progress across a wide range of applications, from robotics to immersive virtual environments.
@@
 By addressing these challenges, the AI community can ensure that advancements in long video reasoning contribute positively to society while mitigating associated risks.
-Appendix C More Demos in Different Reasoning Subsets
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope, as they align with the theoretical and experimental results presented in the paper and provide a clear understanding of the paper’s goals.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The paper acknowledges the limitations of the work performed by the authors.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification:
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: The paper fully discloses all the information needed to reproduce the main experimental results, ensuring that the main claims and conclusions can be independently verified.
+This includes providing relevant details, methodologies, and any necessary parameters or configurations for conducting the experiments.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: The paper will provide open access to the data and code necessary to reproduce the main experimental results.
+It also includes sufficient instructions in the supplemental material on how to faithfully replicate the experiments conducted in the paper.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29.
+Justification: The paper specifies all the training and test details necessary to understand the results.
+This includes information on data splits, hyperparameters, the methodology for selecting hyperparameters, the type of optimizer used, and any other relevant details that are crucial for replicating and comprehending the reported results.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34.
+Justification: The paper provides suitable information about the statistical significance of the experiments.
+This indicates that the authors have appropriately addressed the need for statistical analysis and have reported the relevant measures to support the reliability and significance of their experimental findings.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39.
+Justification: The paper provides sufficient information on the computer resources required to reproduce each experiment.
+This includes details such as the type of compute workers used, the amount of memory required, and the time taken for the execution of the experiments.
+This information allows for accurate replication of the experiments and provides transparency regarding the computational requirements of the study.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: The paper discusses both potential positive and negative societal impacts of the work performed, especially in the resource limited application.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: The paper only use the opensource datasets with legal license.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: The new assets introduced in the paper are well documented and are the documentation provided alongside the assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix
 In this section, we selected examples of four types of reasoning tasks from the benchmark to evaluate and compare the reasoning processes and answers provided by Gemini-1.5-Pro, Video-R1-7B, and LongVILA-R1-7B.
@@
 In contrast, both Video-R1-7B and LongVILA-R1-7B successfully reasoned that the man’s habit of wearing a ring on his left hand is a key indicator, providing accurate answers.
-Figure 9: Detailed data generation process for the LongVideo-Reason dataset, supplementing Figure 4.
-Figure 10: Reward curve of our LongVILA-R1-7B. During training, LongVILA-R1-7B exhibits stable and consistent reward improvements across overall, format, and accuracy dimensions, converging steadily.
 Figure 10: Detailed comparisons in the football game example. The video is available at Link.